### v1

# Leveraging Offline Data in Linear Latent Bandits

Abstract
Sequential decision-making domains such as recommender systems, healthcare and education often have unobserved heterogeneity in the population that can be modeled using latent bandits – a framework where an unobserved latent state determines the model for a trajectory. While the latent bandit framework is compelling, the extent of its generality is unclear. We first address this by establishing a de Finetti theorem for decision processes, and show that every exchangeable and coherent stateless decision process is a latent bandit. The latent bandit framework lends itself particularly well to online learning with offline datasets, a problem of growing interest in sequential decision-making. One can leverage offline latent bandit data to learn a complex model for each latent state, so that an agent can simply learn the latent state online to act optimally. We focus on a linear model for a latent bandit with $d_{A}$-dimensional actions, where the latent states lie in an unknown $d_{K}$-dimensional subspace for $d_{K}\ll d_{A}$. We present SOLD, a novel principled method to learn this subspace from short offline trajectories with guarantees. We then provide two methods to leverage this subspace online: LOCAL-UCB and ProBALL-UCB. We demonstrate that LOCAL-UCB enjoys $\tilde{O}(\min(d_{A}\sqrt{T},d_{K}\sqrt{T}(1+\sqrt{d_{A}T/d_{K}N})))$ regret guarantees, where the effective dimension is lower when the size $N$ of the offline dataset is larger. ProBALL-UCB enjoys a slightly weaker guarantee, but is more practical and computationally efficient. Finally, we establish the efficacy of our methods using experiments on both synthetic data and real-life movie recommendation data from MovieLens.
1 Introduction
Many sequential-decision making problems can be effectively modeled using the bandit framework. This can span domains as diverse as healthcare (Lu et al., 2021), randomized clinical trials (Press, 2009), search and recommendation (Li et al., 2010), distributed networks (Kar et al., 2011), and portfolio design (Brochu et al., 2011). However, in many applications, there often exists unobserved heterogenity in the population, or unobserved context, that influences the distribution of rewards. In Hong et al. (2020), it is shown that this uncertainty can be modeled by a latent bandit (or mixture of bandits). This is a bandit where an unobserved latent state determines the reward model for the trajectory. The latent bandit framework therefore has high practical value.
However, the extent of the generality of this framework is unclear. Given a general decision process, how strong do the conditions need to be to make it a latent bandit? To answer this, we turn to de Finetti’s celebrated theorem, which says that exchangeable sequences of Bernoulli random variables are mixtures of sequences of i.i.d. Bernoulli random variables. This has also been extended to larger classes of stochastic processes. For example, Diaconis and Freedman (1980) show that any random process satisfying recurrence and a weakened notion of exchangeability is a mixture of Markov chains. Inspired by such work, we define a general stateless decision process, and consider a weak notion of exchangeability as well as a natural coherence condition. We show that in the spirit of de Finetti’s theorem, any coherent and exchangeable stateless decision process is a latent bandit.
In light of this equivalence and the empirical relevance of linear contextual bandits (Li et al., 2010; Abbasi-Yadkori et al., 2011; Abbasi-yadkori et al., 2011; Abbasi-Yadkori et al., 2012), we define and explore the setting of linear latent contextual bandits, where a low-dimensional latent state imposes a linear effect on the reward model parameters. When interacting with a latent bandit online, an agent receives rewards based on a single latent state. For example, a patient’s underlying genetic makeup could be the latent parameter that affects their response to treatment. Often, one also has access to large offline datasets with unobserved heterogeneity, like data from randomized controlled trials. Using traditional bandit algorithms in this setting does not leverage the offline data available to the agent. One also cannot simply treat the offline data as coming from a single bandit, since there is significant distribution shift between the mean offline parameter and a single online parameter.
Addressing this, Hong et al. (2020) considers the problem of speeding up online learning using models learnt offline for finitely many latent states. Unlike them, we tackle the problem of using subspaces learnt offline to accelerate online learning, as learning the "latent subspace" spanned by possible reward model parameters is equivalent to learning the infinitely many models in the subspace. We provide a method to learn this subspace from a dataset of trajectories collected under some behavior policy $\pi_{b}$, and seek to use it to speed up online learning. Our subspace learning algorithm is inspired by a long line of work in using spectral methods to learn mixtures of various generative models (Vempala and Wang, 2004; Kong et al., 2020; Chen and Poor, 2022; Kausik et al., 2023). However, each of these methods deals with finitely many latent states, and clusters data under separation assumptions. Clustering is not compatible with continuous latent states, and so, novel ideas are needed to extend these spectral methods to our setting. Finally, the subspace learnt offline is typically an uncertain estimate of the true subspace of reward parameters. Online methods using the subspace should thus account for this uncertainty. We therefore design two such methods with a tractability-tightness tradeoff. With this in mind, we outline our contributions below:
  • We are the first, to our knowledge, to prove a de Finetti theorem for decision processes.
  • We present SOLD, an offline method for learning low-dimensional subspaces of reward parameters with guarantees, generalizing spectral learning methods from discrete to continuous latent states.
  • We present LOCAL-UCB, an online algorithm leveraging subspaces and uncertainty sets estimated offline to sharpen optimism for LinUCB, with $\tilde{O}(\min(d_{A}\sqrt{T},d_{K}\sqrt{T}(1+\sqrt{d_{A}T/d_{K}N})))$ regret.
  • Finally, we present ProBALL-UCB, a practical and computationally efficient online algorithm with a slightly looser regret guarantee. This illustrates a general algorithmic idea for integrating offline subspace estimation into optimistic algorithms.
  • We establish the efficacy of our algorithms outlined above through a simulation study and a demonstration on a real recommendation problem with the MovieLens-1M (Harper and Konstan, 2015) dataset.
Other related work.
The work of Tennenholtz et al. (2021) considers linear contextual bandits with partially observed offline data and fully observable online features, albeit in a very different model of partial information. Additionally, there has been an increasingly large amount of work in offline-to-online transfer learning for reinforcement learning, also known as hybrid RL (Song et al., 2023; Nakamoto et al., 2023; Amortila et al., 2024; Li et al., 2023; Rajeswaran et al., 2017; Hester et al., 2018; Nair et al., 2018, 2020; Ball et al., 2023; Xie et al., 2022; Ren et al., 2024; Wagenmaker and Pacchiano, 2023; Tan and Xu, 2024).
2 A De Finetti Theorem for Decision Processes
We first define a stateless decision process at a high level of generality. [^1] Liu et al. (2023) work with a much more restrictive notion of a generalized bandit and use the original de Finetti theorem in some of their lemmas. See Appendix B.1 for a discussion.
Definition 1.
A stateless decision process (SDP) with action set $\mathcal{A}$ is a probability space $(\Omega,\mathcal{G},\mathbb{P})$ equipped with a family of random maps $\mathcal{F}_{H}:\Omega\to(\mathcal{A}^{H}\to\mathbb{R}^{H})$ for $H\in\mathbb{N}\cup\{\infty\}$.
That is, given a sequence of actions $(a_{1},\dots a_{H})$, an SDP generates a random sequence of rewards $(Y_{1},\dots Y_{H})$. In the light of this interpretation, we will abuse notation to denote by $\mathcal{F}_{H}(a_{1},\dots a_{H})$ the random variable $\omega\mapsto\mathcal{F}_{H}(\omega)(a_{1},\dots a_{H})$. Without any coherence between $\mathcal{F}_{H}$ across $H$, a stateless process can behave arbitrarily for different horizons $H$. We present a natural coherence condition below, essentially requiring that a given action sequence should produce consistent rewards.
Definition 2.
A stateless decision process is said to be coherent if for any $h\leq k\leq H,H^{\prime}\in\mathbb{N}\cup\{\infty\}$ and for any two action sequences $\tau,\tau^{\prime}$ of lengths $H$ and $H^{\prime}$ sharing the same actions $(a_{h},\dots a_{k})$ from index $h$ to $k$, with $\mathcal{F}_{H}(\tau)=(Y_{1},\dots Y_{H})$ and $\mathcal{F}_{H^{\prime}}(\tau^{\prime})=(Y_{1}^{\prime},\dots Y_{H^{\prime}}^{\prime})$, we have $(Y_{h},\dots Y_{k})=(Y_{h}^{\prime},\dots Y_{k}^{\prime})$, viewed as functions of $\Omega$.
It is natural to require equality in value and not just in distribution, since after taking an extra action, the values of past rewards stay the same, not just their distribution. We also give a natural definition for exchangeability of a stateless decision process – namely that exchanging any two rewards should lead to the distribution obtained by exchanging the corresponding actions.
Definition 3.
A stateless decision process is said to be exchangeable if for any permutation $\pi:[H]\to[H]$ and $\mathcal{F}_{H}(a_{1},\dots a_{H})=(Y_{1},\dots Y_{H})$, we have $\mathcal{F}_{h}(a_{\pi(1)},\dots a_{\pi(H)})\sim(Y_{\pi(1)},\dots Y_{\pi(H)})$.
Finally, a latent bandit is an SDP that behaves like a bandit conditioned on a random latent state $F$ that determines the reward distribution. As $F$ determines a distribution, it is a random measure-valued function on $\mathcal{A}$.
Definition 4.
A latent bandit is a stateless decision process equipped with a random measure-valued function $F:\Omega\to(\mathcal{A}\to\mathcal{P}(\mathbb{R}))$ so that for any $H$ and action sequence $(a_{1},\dots a_{H})$, the rewards $(Y_{1},\dots Y_{H}):=\mathcal{F}_{H}(a_{1},\dots a_{H})$ are independent conditioned on $F$. Moreover, the conditional distribution $\mathcal{L}[Y_{h}\mid F]=F(a_{h})$ for all $h\leq H$ [^2] We abuse notation twice here. First, we write $F(a_{h}):=(\omega\mapsto F(\omega)(a_{h}))$. Second, as the regular conditional distribution $\mathcal{L}[Y_{h}\mid F]$ is a kernel that maps from $\Omega\times\mathcal{B}\to\mathbb{R}$, we view $F(a_{h})$ as its curried map $(\omega,B)\mapsto F(a_{h})(\omega)(B)$. A discussion of issues like measurability and well-definedness is in Appendix B.2.
While exchangeability and coherence are reasonable conditions on an SDP and are clearly satisfied by latent bandits, it is a-priori unclear if they are sufficient to ensure that the SDP is a latent bandit. In particular, since only exchanging rewards from the same action preserves the distribution, standard de Finetti proof ideas do not immediately apply. After all, it is possible that an SDP could be cleverly designed to choose rewards adaptively across time and satisfy these properties. Reassuringly, no such counterexamples exist and we have the following theorem.
Theorem 1 (De Finetti Theorem for Stateless Decision Processes).
Every exchangeable and coherent stateless decision process is a latent bandit.
We show in Lemma 2 that coherence is not a consequence of exchangeability – it is a necessary condition for being a latent bandit. Finally, we consider contexts and define contextual decision processes in a manner agnostic to context transitions. That is, the process only carries the data of how rewards are generated from given sequences of contexts and actions, while the context transitions themselves may be generated by a different process.
Definition 5.
A transition-agnostic contextual decision process (TACDP) with action set $\mathcal{A}$ and context set $\mathcal{X}$ is a probability space $(\Omega,\mathcal{G},\mathbb{P})$ equipped with a family of random maps $\mathcal{F}_{H}:\Omega\to((\mathcal{X}\times\mathcal{A})^{H}\to\mathbb{R}^{H})$ for $H\in\mathbb{N}$ as well as a map $\mathcal{F}_{\infty}:\Omega\to((\mathcal{X}\times\mathcal{A})^{\mathbb{N}}\to\mathbb{R}^{\mathbb{N}})$.
It is natural to isolate away context dynamics for processes like stochastic contextual bandits, in which actions do not affect the context transitions – in contrast to transition-aware definitions in Jiang et al. (2016). We can analogously define coherence and exchangeability for TACDPs and define latent contextual bandits by simply replacing $\mathcal{A}$ with $\mathcal{X}\times\mathcal{A}$ in the definitions above. We then show an analogous de Finetti theorem, as a corollary of our proof of Theorem 1. See Appendix B.2 for more details.
3 Linear Latent Contextual Bandits
Latent bandits are a powerful and general framework for encoding uncertainty in reward models. However, this generality is both a blessing and a curse. Without further assumptions on the structure of the latent state $F$, one cannot hope to estimate it. We therefore focus on a special case here. We consider the natural generalization of the linear contextual bandit to the latent bandit setting, where we impose a linear structure on the effect of latent states. We further justify this by noting that in most application domains, it is reasonable to assume that a parsimonious, low-dimensional latent state affects the reward distribution. This motivates the following definition.
Definition 6.
A linear latent contextual bandit is a TACDP equipped with a feature map for context-action pairs $\phi:\mathcal{X}\times\mathcal{A}\to\mathbb{R}^{d_{A}}$, a latent random variable $\bm{\theta}\in\mathbb{R}^{d_{K}}$ with distribution $\mathcal{D}_{\theta}$ and a map $\textbf{U}_{\star}:\mathbb{R}^{d_{K}}\to\mathbb{R}^{d_{A}}$ such that for any $H$ and context-action sequence $((x_{1},a_{1}),\dots(x_{H},a_{H}))$, the rewards $(Y_{1},\dots Y_{H})$ are independent conditioned on $\bm{\theta}$. Moreover, $Y_{h}\mid\bm{\theta}\sim\phi(x_{h},a_{h})^{\top}\bm{\beta}+\epsilon$, where $\epsilon$ is subgaussian noise independent of all actions and all other observations, and $\bm{\beta}=\textbf{U}_{\star}\bm{\theta}$.
In other words, this is a latent contextual bandit where the random measure-valued function $F:\Omega\to((\mathcal{X}\times\mathcal{A})\to\mathcal{P}(\mathbb{R}))$ is defined by setting $F(\omega)(x,a)$ to be the distribution given by $\phi(x,a)^{\top}\textbf{U}_{\star}\bm{\theta}(\omega)+\epsilon$, for any $\omega\in\Omega$ and $(x,a)\in(\mathcal{X}\times\mathcal{A})$. Further, we note that WLOG $\textbf{U}_{\star}$ is unitary. This is because for any invertible map $A:\mathbb{R}^{d_{K}}\to\mathbb{R}^{d_{K}}$, the observation distribution does not change upon replacing $\bm{\theta}$ with $A\bm{\theta}$ and replacing $\textbf{U}_{\star}$ with $\textbf{U}_{\star}A^{-1}$. One can see this as a generalization of the fact that with finitely many latent states, permuting the labels and rewards together does not change observations.
Let us now assume that we have access to a dataset $\mathcal{D}_{\operatorname{off}}$ of $N$ short trajectories $\tau_{n}=((x_{n,1},a_{n,1},r_{n,1}),...,$ $(x_{n,H},a_{n,H},r_{n,1}))$ of length $H$, collected by some behavior policy $\pi_{b}$. In online deployment, a single latent label $\bm{\theta}_{\star}$ is chosen and rewards are generated using $\bm{\beta}_{\star}=\textbf{U}_{\star}\bm{\theta}_{\star}$. At each timestep $t$, an agent observes contexts $x_{t}$ and uses both the offline data and the online data at time $t$ to execute a policy $\pi_{t}$. Define the optimal action at time $t$ by $a^{\star}_{t}:=\max_{a}\phi(x_{t},a_{t})^{\top}\bm{\beta}_{\star}$ We tackle the problem of minimizing the frequentist regret in linear latent contextual bandits, given by
$\operatorname{Reg}_{T}:=\textstyle{\sum_{t=1}^{T}\phi(x_{t},a^{\star}_{t})^{\top}\bm{\beta}_{\star}-\mathbb{E}_{a\sim\pi_{t}}[\phi(x_{t},a)^{\top}\bm{\beta}_{\star}].}$
For example, in medical applications, data from short randomized controlled trials can be used to help an agent suggest treatment decisions for a new patient online. In this case, we would like the algorithm to administer the correct treatments for each patient. This means that the frequentist regret is the relevant performance metric here, and not the Bayesian regret over some prior. Additionally, any worst-case bound on the frequentist regret is a bound on the Bayesian regret for arbitrary priors.
Challenges with latent bandits. Despite the linear assumption, and the dimension reduction obtained in the common case when $d_{K}\ll d_{A}$, significant challenges remain. First, the value of the latent state $\bm{\theta}$ and the map $\textbf{U}_{\star}$ are both unknown a priori, making it hard to leverage the low-dimensional structure of the problem. Second, a good choice of dimension $d_{K}$ is itself unknown a priori, and must be determined from data in a principled manner. In the following section, we will provide a method to estimate latent subspaces from offline data that allows us to overcome these challenges.
4 Estimating Latent Subspaces from Offline Data
Although we do not have access to the values of the latent states $\bm{\theta}$ or to the map $\textbf{U}_{\star}$, we can still extract useful information from data. To that effect, recall that we have access to a dataset $\mathcal{D}_{\operatorname{off}}$ of $N$ trajectories $\tau_{n}=((x_{n,h},a_{n,h},r_{n,h}))_{h=1}^{H}$ of length $H$, collected by some behavior policy $\pi_{b}$.
How can offline data help us in online deployment?
To minimize the regret, one must learn $\bm{\beta}_{\star}$ online. However, it is much easier to search among all $\bm{\theta}\in\mathbb{R}^{d_{K}}$ than to search among all possible $\bm{\beta}\in\mathbb{R}^{d_{A}}$ since typically, $d_{K}\ll d_{A}$. So, it will help to learn some $\hat{\textbf{U}}^{\top}\approx\textbf{U}_{*}^{\top}:\mathbb{R}^{d_{A}}\to\mathbb{R}^{d_{K}}$ offline so that for any estimate $\hat{\bm{\beta}}_{t}$ of $\bm{\beta}_{\star}$, $\hat{\textbf{U}}^{\top}\hat{\bm{\beta}}_{t}$ is an estimate of $\hat{\bm{\theta}}_{t}\in\mathbb{R}^{d_{K}}$. This amounts to learning a subspace of the feature space from logged bandit data. We therefore provide a method for Subspace estimation from Offline Latent bandit Data (SOLD), presented as Algorithm 1.
Algorithm 1 Subspace estimation from Offline Latent bandit Data (SOLD)
1:  Input: Dataset $\mathcal{D}_{\operatorname{off}}$ of collected trajectories $\tau_{n}=((x_{n,1},a_{n,1},r_{n,1}),...,(x_{n,H},a_{n,H},r_{n,1}))$ under a behavior policy $\pi_{b}$, dimension of latent subspace $d_{K}$.
2:  Divide each $\tau_{n}$ into odd and even steps, giving trajectory halves $\tau_{n,1}$ and $\tau_{n,2}$.
3:  Estimate reward parameters $\hat{\bm{\beta}}_{n,i}\leftarrow\textbf{V}_{n,i}^{-1}\textbf{b}_{n,i}$, where $\textbf{V}_{n,i}=\mu\textbf{I}+\sum_{(x,a,r)\in\tau_{n,i}}\phi(x,a)\phi(x,a)^{\top}$ and $\textbf{b}_{n,i}=\sum_{(x,a,r)\in\tau_{n,i}}\phi(x,a)r$ for $i=1,2$.
4:  Compute $\textbf{M}_{n}=\frac{1}{2}(\hat{\bm{\beta}}_{n,1}\hat{\bm{\beta}}_{n,2}^{\top}+\hat{\bm{\beta}}_{n,2}\hat{\bm{\beta}}_{n,1}^{\top})$ and compute $\overline{\textbf{M}}_{N}:=\frac{1}{N}\sum_{n=1}^{N}\textbf{M}_{n}$.
5:  Compute matrices $\overline{\textbf{D}}_{N,i}:=\frac{1}{N}\sum_{n=1}^{N}(I-\mu\textbf{V}_{n,i}^{-1})$ for $i=1,2$.
6:  Obtain the top $d_{K}$ eigenvectors of $\overline{\textbf{D}}_{N,1}^{-1}{\overline{\textbf{M}}_{N}}{\overline{\textbf{D}}_{N,2}^{-1}}$, denoting them by $\hat{\textbf{U}}$.
7:  return  Projection matrix $\hat{\textbf{U}}\hat{\textbf{U}}^{\top}$, $\Delta_{\operatorname{off}}$ as in Theorem 2
On trajectory splitting and corrections.
We split each trajectory to obtain two independent estimators $\hat{\bm{\beta}}_{n,1},\hat{\bm{\beta}}_{n,2}$ of $\bm{\beta}$, compute the outer products $\hat{\bm{\beta}}_{n,1}^{\top}\hat{\bm{\beta}}_{n,2}$ for each trajectory, and try to obtain the top $d_{K}$ eigenvectors of the mean outer product across trajectories. We want the mean outer product to concentrate to $\mathbb{E}[\bm{\beta}\bm{\beta}^{\top}]=\textbf{U}_{\star}\mathbb{E}[\bm{\theta}\bm{\theta}^{\top}]\textbf{U}_{\star}^{\top}$, which has the same $d_{K}$-dimensional span as $\textbf{U}_{\star}$.
However, there is a wrinkle here. We cannot simply take the top $d_{K}$ eigenvectors of the mean outer product $\overline{\textbf{M}}_{N}$. Note that $\mathbb{E}[\overline{\textbf{M}}_{N}]=\mathbb{E}[\textbf{D}_{n,1}\bm{\beta}_{n}\bm{\beta}_{n}^{\top}\textbf{D}_{n,2}]=\mathbb{E}[\textbf{D}_{n,1}\textbf{U}_{\star}\bm{\theta}_{n}\bm{\theta}_{n}^{\top}\textbf{U}_{\star}^{\top}\textbf{D}_{n,2}]$. To separate $\mathbb{E}[\bm{\beta}_{n}\bm{\beta}_{n}^{\top}]$ from this, we need $\textbf{D}_{n,1},\bm{\beta}_{n},\textbf{D}_{n,2}$ to be independent. If $\pi_{b}$ does not use $\bm{\theta}$ and contexts are generated independently of each other and of $\bm{\theta}$, then this is satisfied. In fact, we show in the lemma below that if any of these three conditions is violated, then it is in fact impossible to determine the latent subspace $\textbf{U}^{\star}$ using any method, even with infinitely many infinitely long trajectories.
Lemma 1 (Contexts, $\bm{\theta}$, and $\pi_{b}$ cannot be dependent).
For each of these conditions:
  1. 1.
    Contexts in a trajectory are dependent but do not depend on $\bm{\theta}$, and $\pi_{b}$ also does not use $\bm{\theta}$,
  2. 2.
    Contexts are generated independently using $\bm{\theta}$, while $\pi_{b}$ does not use $\bm{\theta}$,
  3. 3.
    Contexts are generated independently without using $\bm{\theta}$, while $\pi_{b}$ uses $\bm{\theta}$,
there exists a pair of linear latent contextual bandits with orthogonal latent subspaces satisfying the condition, and a behavior policy $\pi_{b}$ so that the offline data distributions are indistinguishable and cover all $(x,a)$ pairs with probability at least $1/4$. Since the latent subspaces are orthogonal, an action that gives the maximum reward on one latent bandit gives reward 00 on the other.
To estimate the latent subspace, one is thus forced to make the following assumption.
Assumption 1 (Unconfounded Offline Actions).
The offline behavior policy $\pi_{b}$ does not use $\bm{\theta}$ to choose actions, and contexts $x_{n,h}$ are stochastic and generated independently of each other and of $\bm{\theta}$.
As an example, this assumption is satisfied in practice when the offline data comes from randomized controlled trials or A/B testing, which are common sources of offline datasets. [^3] Even when the assumption is not satisfied in practice, Algorithm 1 can learn a good subspace whenever $\overline{\textbf{D}}_{N,1}^{-1}{\overline{\textbf{M}}_{N}}{\overline{\textbf{D}}_{N,2}^{-1}}$ has eigenspace close to the span of $\textbf{U}_{\star}$, e.g. when high-reward actions contribute heavily to $\textbf{D}_{n,i}$. Now, let the covariance matrix of $\bm{\theta}$ be $\Lambda$ and let its mean be $\mu_{\theta}$. Then we have that $\mathbb{E}[\overline{\textbf{M}}_{N}]=\mathbb{E}[\textbf{D}_{n,1}\bm{\beta}_{n}\bm{\beta}_{n}^{\top}\textbf{D}_{n,2}]=\mathbb{E}[\textbf{D}_{n,1}]\textbf{U}_{\star}(\Lambda+\mu_{\theta}\mu_{\theta}^{\top})\textbf{U}_{\star}^{\top}\mathbb{E}[\textbf{D}_{n,2}]$. So, we still cannot merely consider the top $d_{K}$ eigenvectors of $\overline{\textbf{M}}_{N}$ without accounting for $\textbf{D}_{n,1}$. Intuitively, $\textbf{D}_{n,1}$ captures the fact that we cannot estimate rewards of actions that we have not taken in a short trajectory. We therefore construct correction matrices $\overline{\textbf{D}}_{N,i}$ and use them to "neutralize" the effect of unseen actions from the short trajectories. In particular, $\overline{\textbf{D}}_{N,1}^{-1}{\overline{\textbf{M}}_{N}}{\overline{\textbf{D}}_{N,2}^{-1}}$ is an estimator for $\textbf{U}_{\star}(\Lambda+\mu_{\theta}\mu_{\theta}^{\top})\textbf{U}_{\star}^{\top}$. Crucially, this allows us to aggregate information across many trajectories to overcome the challenge of learning from short trajectories. We can now take the top $d_{K}$ eigenvectors of $\overline{\textbf{D}}_{N,1}^{-1}{\overline{\textbf{M}}_{N}}{\overline{\textbf{D}}_{N,2}^{-1}}$ to estimate the subspace determined by $\textbf{U}_{\star}$. However, to give guarantees for our method, we must first make a coverage assumption:
Assumption 2 (Boundedness and Coverage).
Rewards $|r_{n,h}|\leq R$ for all $n,h$, $\|\phi(x,a)\|_{2}\leq 1$ and $\|\bm{\beta}\|_{2}\leq R$. Also, $\lambda_{A}:=\min_{i=1,2}\lambda_{\min}(\mathbb{E}[\textbf{D}_{n,i}])>0$ and $\lambda_{\theta}:=\frac{1}{R^{2}}\lambda_{\min}(\Lambda)>0$.
Intuitively, $\lambda_{A}$ measures coverage along actions, while $\lambda_{\theta}$ measures coverage along latent states $\bm{\theta}$. Both must be non-zero to expect satisfactory estimation of the subspace. We can then use confidence bounds for $\overline{\textbf{M}}_{N}$ and $\overline{\textbf{D}}_{N,i}$ to give a data-dependent confidence bounds $\Delta_{\operatorname{off}}$ for the projection matrix $\hat{\textbf{U}}\hat{\textbf{U}}^{\top}$, as in Theorem 2 below. In one instantiation, Propositions 1 and 2 in Appendix C derive simple data-dependent bounds for $\overline{\textbf{M}}_{N}$ and $\overline{\textbf{D}}_{N,i}$ respectively. Under this choice, we can control the growth of $\Delta_{\operatorname{off}}$ in terms of the unknown problem parameters, shown at the end of Theorem 2.
Theorem 2 (Computing and Bounding $\Delta_{\operatorname{off}}$).
Let $\|\overline{\textbf{M}}_{N}-\mathbb{E}[\textbf{M}_{1}]\|_{2}\leq\Delta_{M}$ and $\|\overline{\textbf{D}}_{N,i}-\mathbb{E}[\textbf{D}_{n,i}]\|_{2}\leq\Delta_{D}$ for $i=1,2$ with probability $31-\delta/$ 3 each. Then, with probability $1-\delta$, $\|\hat{\textbf{U}}\hat{\textbf{U}}^{\top}-\textbf{U}_{\star}\textbf{U}_{\star}^{\top}\|_{2}\leq\Delta_{\operatorname{off}}$, where for $B_{D}=\|\overline{\textbf{D}}_{N}^{-1}\|_{2}$ and $\hat{\lambda}:=\lambda_{d_{K}}(\overline{\textbf{M}}_{N})-\lambda_{d_{K}+1}(\overline{\textbf{M}}_{N})$,
$\Delta_{\operatorname{off}}=\textstyle{\frac{2\sqrt{2d_{K}}}{\hat{\lambda}}\left(\left(\frac{B_{D}^{3}(2-B_{D}\Delta_{D})}{(1-B_{D}\Delta_{D})^{2}}\right)(R^{2}+\Delta_{M})\Delta_{D}+\left(\frac{B_{D}}{1-B_{D}\Delta_{D}}\right)^{2}\Delta_{M}\right)}$
Obtaining $\Delta_{M}$ and $\Delta_{D}$ from Propositions 1 and 2, $\Delta_{\operatorname{off}}=\widetilde{O}({\frac{1}{\lambda_{\theta}\lambda_{A}^{3}}N^{-1/2}}\sqrt{d_{K}d_{A}\log(d_{A}/\delta)})$.
Estimating $d_{K}$ from offline data.
Since our estimator $\overline{\textbf{D}}_{N,1}^{-1}{\overline{\textbf{M}}_{N}}{\overline{\textbf{D}}_{N,2}^{-1}}$ is approximately rank-$d_{K}$, the number of significant eigenvalues of the estimator is a principled heuristic for determining $d_{K}$.
Insufficiency of PCA and PMF for subspace estimation.
Naively performing PCA on the raw rewards or on single reward estimates $\hat{\beta}_{n}$ can lead to erroneous subspaces – as while the PCA target is linear-algebraically similar to $\overline{\textbf{M}}_{N}$, it is statistically different. The PCA target (e.g. $\mathbb{E}[\hat{\beta}_{n}\hat{\beta}_{n}^{\top}]$) is typically full rank due to the variance of the per-reward noise $\epsilon$. On the other hand, PMF (Mnih and Salakhutdinov, 2007b) offers neither confidence bounds on the estimated subspace, nor a principled method for determining $d_{K}$.
5 Using Offline Data to Speed Up Online Learning with LOCAL-UCB
Here, we motivate and describe LOCAL-UCB, a natural algorithm that accelerates LinUCB with offline data. The core idea is sharpening optimism by being optimistic over the intersection of two confidence sets – one obtained using offline data and another purely from online data.
To geometrically motivate our update rule, say $t$ online interactions have been completed, yielding a dataset $\mathcal{D}_{t}$ of $(x_{s},a_{s},r_{s})$ tuples. From Theorem 2, SOLD gives an estimate and a confidence set for the projection matrix $\textbf{U}_{\star}\textbf{U}_{\star}^{\top}$. Let’s call the set $\mathcal{C}_{U,\operatorname{off}}$. For each projection $\textbf{U}\textbf{U}^{\top}\in\mathcal{C}_{U,\operatorname{off}}$, we have a $d_{K}$-dimensional confidence ellipsoid for $\bm{\beta}$ within the corresponding subspace. The union of all such $d_{K}$-dimensional confidence ellipsoids creates one confidence set for $\bm{\beta}$, coming from offline uncertainty. Let’s call it $\mathcal{C}^{t}(\beta,\operatorname{off})$. On the other hand, we also have the purely online $d_{A}$-dimensional confidence ellipsoid $\mathcal{C}^{t}(\beta,\operatorname{on})$. We can thus sharpen our optimism by being optimistic over the smaller intersection $\mathcal{C}^{t}(\beta,\operatorname{off})\cap\mathcal{C}^{t}(\beta,\operatorname{on})$ instead of just over $\mathcal{C}^{t}(\beta,\operatorname{on})$ – thereby performing "more localized" optimism than LinUCB. This is the idea behind Algorithm 2, which we thus call LOCAL-UCB. Figure 1 illustrates different ways in which taking this intersection can affect optimism.
We formalize this intuition in Algorithm 2 by formulating the sharpened optimism as an optimization problem in Line 4. The first constraint respresents the low dimensional confidence ellipsoids in the subspace spanned by a given U, while the third constraint lets U range over all of $\mathcal{C}_{U,\operatorname{off}}$. The second constraint merely represents the usual high dimensional ellipsoid.
Algorithm 2 Latent Offline subspace Constraints for Accelerating Linear UCB (LOCAL-UCB)
1:  Input: Projection matrix $\hat{\textbf{U}}\hat{\textbf{U}}^{\top}$ and confidence bound $\Delta_{\operatorname{off}}$ from an offline uncertainty-aware method, e.g. SOLD.
2:  Initialize $\textbf{V}_{1}\leftarrow\textbf{I}_{d_{A}}$, $\textbf{b}_{1}\leftarrow 0$, $\alpha_{t}$
3:  for $t=1,\dots T$ do
4:     Play action $a_{t}$ and receive reward $r_{t}$ according to:
$\textstyle{a_{t},\widetilde{\bm{\beta}}_{t},\widetilde{\textbf{U}}_{t}\leftarrow\arg\max_{a,\bm{\beta},\textbf{U}}\phi(x_{t},a)^{\top}\bm{\beta}}\text{ such that }$
$\textstyle{\hat{\bm{\beta}}_{1,t}\leftarrow\textbf{U}(\textbf{U}^{T}\textbf{V}_{t}\textbf{U})^{-1}\textbf{U}^{\top}\textbf{b}_{t},\ \|\textbf{U}^{\top}(\bm{\beta}-\hat{\bm{\beta}}_{1,t})\|_{(\textbf{U}^{\top}\textbf{V}_{t}\textbf{U})^{-1}}\leq\alpha_{1,t}}$
$\textstyle{\hat{\bm{\beta}}_{2,t}\leftarrow\textbf{V}_{t}^{-1}\textbf{b}_{t},\|(\bm{\beta}-\hat{\bm{\beta}}_{2,t})\|_{\textbf{V}_{t}^{-1}}\leq\alpha_{2,t}}$
$\textstyle{\|\bm{\beta}\|_{2}\leq R,\textbf{U}^{\top}\textbf{U}=\textbf{I}_{d_{K}},\textbf{U}\textbf{U}^{\top}\bm{\beta}=\bm{\beta},\ \|\hat{\textbf{U}}^{\top}\textbf{U}\|_{F}\geq\sqrt{d_{K}-\Delta_{\operatorname{off}}^{2}/2}}$
5:     Compute $\textbf{b}_{t+1}\leftarrow\textbf{b}_{t}+\phi(x_{t},a)r_{t}$, $\textbf{V}_{t+1}\leftarrow\textbf{V}_{t}+\phi(x_{t},a)\phi(x_{t},a)^{\top}$, $\alpha_{t+1}$
6:  end for
We provide the following guarantee for LOCAL-UCB. Notice that our guarantee shows that for enough offline data with $N\gg T$, the effective dimension of the problem is $d_{K}$. It increases to $d_{A}$ as $T$ gets closer to $N$. The quality of the offline data is reflected in the coverage constants $\lambda_{D}$ and $\lambda_{\min}$.
Theorem 3 (LOCAL-UCB Regret).
Under Assumptions 1 and 2, if $\alpha_{1,t}=R\sqrt{\mu}+CR\sqrt{d_{K}\log(2T/\delta)}$ and $\alpha_{2,t}=R\sqrt{\mu}+CR\sqrt{d_{A}\log(2T/\delta)}$ for a suitable universal constant $C$, with probability at least $1-\delta$, LOCAL-UCB has regret
$\operatorname{Reg}_{T}$ $=O\textstyle{\left(\min\left(Rd_{A}\sqrt{T},Rd_{K}\sqrt{T}\left(1+\frac{1}{\lambda_{\theta}\lambda_{A}^{3}}\sqrt{\frac{Td_{A}}{d_{K}N}}\right)\right)\right)}.$
However, the subspace constraint $\|\hat{\textbf{U}}^{\top}\textbf{U}\|_{F}\geq\sqrt{d_{K}-\Delta_{\operatorname{off}}^{2}/2}$ is nonconvex. In fact, we lower bound a convex function in the constraint. This essentially searches over a complicated star-shaped set in the space of all U. As such, LOCAL-UCB is unfortunately computationally inefficient.
Figure 1: Left: Geometric interpretations of LOCAL-UCB. Showing $\mathcal{C}^{t}(\bm{\beta},\operatorname{on})\cap\mathcal{C}^{t}(\bm{\beta},\operatorname{off})$ in green for three timepoints $t=t_{1},t_{2},t_{3}$. The dotted lines delineate the subspace confidence set. Right: Geometric interpretation of ProBALL-UCB. $\mathcal{C}^{t_{1}}(\bm{\beta},\operatorname{on})\not\subset\widetilde{\mathcal{C}}^{t_{1}}(\bm{\beta},\operatorname{off})$, so we continue to use projections; but by time $t_{2},$ $\mathcal{C}^{t_{2}}(\bm{\beta},\operatorname{on})\subset\widetilde{\mathcal{C}}^{t_{2}}(\bm{\beta},\operatorname{off})$, so we stop using projections.
6 More Practical Optimism with ProBALL-UCB
We note that while LOCAL-UCB is a natural algorithm with promising guarantees, it is not computationally efficient. We address this issue by introducing ProBALL-UCB (Algorithm 3), a practical and computationally efficient algorithm for accelerating online learning in linear latent bandits with offline data. We make three simplifications compared to LOCAL-UCB. Namely, (1) this has cruder confidence sets, (2) we wait for one confidence set to become a subset of the other instead of considering complex intersections, and (3) we use a computable proxy for the above condition instead of explicitly checking it. Surprisingly, this leads to only a minor degradation in provable guarantees.
To geometrically motivate our update rule like in Section 5, we consider the $d_{K}$-dimensional confidence ellipsoid $\mathcal{C}^{t}(\bm{\beta},\hat{\textbf{U}})$ of $\bm{\beta}$ inside subspace $\hat{\textbf{U}}^{\top}$. First, since $\|\hat{\textbf{U}}\hat{\textbf{U}}^{\top}-\textbf{U}_{\star}\textbf{U}_{\star}^{\top}\|_{2}\leq\Delta_{\operatorname{off}}$, we can show that a crude bound for $\phi(x,a)^{\top}\bm{\beta}_{\star}$ is $\max_{\bm{\beta}\in\mathcal{C}^{t}(\bm{\beta},\hat{\textbf{U}})}\phi(x,a)^{\top}\bm{\beta}+\Delta_{\operatorname{off}}$. This corresponds to optimism over a crude confidence ball $\widetilde{\mathcal{C}}^{t}(\beta,\operatorname{off})$ of radius $\Delta_{\operatorname{off}}$ around $\mathcal{C}^{t}(\beta,\hat{\textbf{U}})$. Second, instead of being optimistic over the intersection of this ball with our usual $d_{A}$-dimensional confidence ellipsoid $\mathcal{C}^{t}(\beta,\operatorname{on})$, we merely wait for $\mathcal{C}^{t}(\beta,\operatorname{on})$ to be inside $\widetilde{\mathcal{C}}^{t}(\beta,\operatorname{off})$. Finally, instead of actually checking whether $\mathcal{C}^{t}(\beta,\operatorname{on})$ is inside $\widetilde{\mathcal{C}}^{t}(\beta,\operatorname{off})$, we use a proxy condition that roughly demands $\Delta_{\operatorname{off}}T>d_{A}\sqrt{T}$. Here, $d_{A}\sqrt{T}$ represents the "cumulative size" of the ellipsoid across timesteps $t$, while $\Delta_{\operatorname{off}}T$ represents the "cumulative size" of the ball. This is illustrated in Figure 1.
We note that the idea behind ProBALL-UCB is fairly general and can be used to design adaptations of other algorithms as well. Broadly, we project to a subspace, perform optimism there, and track the cumulative error for using this subspace. Once the cumulative error of using the subspace ($\approx\Delta_{\operatorname{off}}T$ in our case) exceeds the cumulative error of not using the subspace ($\approx d_{A}\sqrt{T}$), we stop using the subspace. Although ProBALL-UCB instantiates this framework with LinUCB, the same idea can be immediately applied to other bonus-based algorithms like SupLinUCB.
Algorithm 3 Projection and Bonuses for Accelerating Latent bandit Linear UCB (ProBALL-UCB)
1:  Input: Projection matrix $\hat{\textbf{U}}\hat{\textbf{U}}^{\top}$, confidence bound $\Delta_{\operatorname{off}}$. Hyperparameters $\alpha_{1,t},\alpha_{2,t},\tau,\tau^{\prime}$.
2:  Initialize $\textbf{V}_{1}\leftarrow I$, $\textbf{b}_{1}\leftarrow 0$, $\textbf{C}_{t}\leftarrow 0$
3:  for $t=1,\dots T$ do
4:     if $\Delta_{\operatorname{off}}\tau\sqrt{t}+\Delta_{\operatorname{off}}\tau^{\prime}\sqrt{{d_{K}}\textstyle{\sum}_{s=1}^{t}\kappa_{s}^{2}/t}\leq d_{A}$ then
5:        Compute $\hat{\bm{\beta}}_{1,t}\leftarrow\hat{\textbf{U}}(\hat{\textbf{U}}^{\top}\textbf{V}_{t}\hat{\textbf{U}})^{-1}\hat{\textbf{U}}^{\top}\textbf{b}_{t}$
6:        Play $a_{t}\leftarrow\arg\max_{a}\phi(x_{t},a)^{\top}\hat{\textbf{U}}\hat{\textbf{U}}^{\top}\hat{\bm{\beta}}_{1,t}+\alpha_{1,t}\|\phi(x_{t},a)^{\top}\hat{\textbf{U}}\|_{(\hat{\textbf{U}}^{\top}\textbf{V}_{t}\hat{\textbf{U}})^{-1}}$
7:     else
8:        Compute $\hat{\bm{\beta}}_{2,t}\leftarrow\textbf{V}_{t}^{-1}\textbf{b}_{t}$
9:        Play $a_{t}\leftarrow\arg\max_{a}\phi(x_{t},a)^{\top}\hat{\bm{\beta}}_{2,t}+\alpha_{2,t}\left\|\phi(x_{t},a)\right\|_{\textbf{V}_{t}^{-1}}$
10:     end if
11:     Observe reward $r_{t}$ and update $\textbf{b}_{t+1}\leftarrow\textbf{b}_{t}+\phi(x_{t},a)r_{t}$, $\textbf{V}_{t+1}\leftarrow\textbf{V}_{t}+\phi(x_{t},a)\phi(x_{t},a)^{\top}$
12:     Update $\textbf{C}_{t+1}\leftarrow\textbf{C}_{t}+\hat{\textbf{U}}^{\top}\phi(x_{t},a_{t})\phi(x_{t},a_{t})^{\top}$, $\kappa_{t+1}\leftarrow\|\textbf{C}_{t+1}\|_{(\hat{\textbf{U}}^{\top}\textbf{V}_{t+1}\hat{\textbf{U}})^{-1}}$
13:  end for
As a final note before the regret bound, there is a technical challenge with analyzing ProBALL-UCB. Since $\hat{\bm{\beta}}_{1,t}$ lies in $\hat{\textbf{U}}$ but $\bm{\beta}_{\star}$ might not, the $d_{K}$-dimensional confidence ellipsoid bound no longer applies. We therefore prove our own confidence ellipsoid bound in Lemma 6, to bypass this issue.
Theorem 4 (Regret for ProBALL-UCB).
Let $\alpha_{1,t}=R\sqrt{\mu}+\tau^{\prime}R\Delta_{\operatorname{off}}\kappa_{t}+CR\sqrt{d_{K}\log(T/\delta)}$ and let $\alpha_{2,t}=R\sqrt{\mu}+CR\sqrt{d_{A}\log(T/\delta)}$. Let $S$ be the first timestep when Algorithm 3 does not play Line 6 and let $S=T$ if no such timestep exists. For $\tau=\tau^{\prime}=1$ we have that
$\operatorname{Reg}_{T}$ $=\textstyle{\widetilde{O}\left(\min\left(Rd_{A}\sqrt{T},Rd_{K}\sqrt{T}\left(1+\frac{1}{\lambda_{A}^{3}\lambda_{\theta}}\left(\sqrt{\frac{d_{A}T}{d_{K}N}}+\sqrt{\frac{d_{A}}{SN}\sum_{t=1}^{S}\kappa_{t}^{2}}\right)\right)\right)\right)}.$
In the worst case, $\kappa_{t}=O(t)$ and so $\frac{1}{S}\sum_{t=1}^{S}\kappa_{t}^{2}=O(T^{2})$, but if all features $\phi(x_{t},a_{t})$ lie in the span of $\hat{\textbf{U}}$ for $t\leq S$, then $\frac{1}{S}\sum_{t=1}^{S}\kappa_{t}^{2}=O(T)$.
While the regret bound looks weaker in the worst case, we emphasize that the "good case" in Theorem 4 is quite common. As an illustrative example, if the feature set $\mathcal{F}_{t}=\{\phi(x_{t},a)\mid a\in\mathcal{A}\}$ is an $\ell_{2}$ ball, then the maximization problem in Step 6 will always choose $a_{t}$ with $\phi(x_{t},a_{t})$ in the span of $\hat{\textbf{U}}$. This can also approximately hold if the features are roughly isotropic or close to the span of $\hat{\textbf{U}}$.
7 Experiments
We now establish the practical efficacy of SOLD (Algorithm 1) and ProBALL-UCB (Algorithm 3) for linear latent contextual bandits through a series of numerical experiments. We perform a simulation study and a demonstration using real-life data. [^4] All experiments were run on a single computer with an Intel i9-13900k CPU, 128GB of RAM, and a NVIDIA RTX 3090 GPU, in no more than an hour in total. We obtain confidence bounds $\Delta_{\operatorname{off}}$ using three different concentration inequalities – (1) Hoeffding as in Proposition 2 (H-ProBALL), (2) empirical Bernstein as in Proposition 1 (E-ProBALL), and (3) confidence intervals generated by the martingale Bernstein concentration inequalities of Waudby-Smith and Ramdas (2023); Wang and Ramdas (2024) (M-ProBALL). We use a simpler expression for $\Delta_{\operatorname{off}}$, set $\tau^{\prime}=0$, and vary the hyperparameter $\tau$ to adjust for overly conservative $\Delta_{\operatorname{off}}$ [^5] Namely, we set $\Delta_{D}=0$ in $\Delta_{\operatorname{off}}$. Also, Lemma 6 and some thought reveal that choosing $\tau^{\prime}=0$ recovers the ”good” version of ProBALL-UCB guarantees, if features are isotropic enough..
Simulation study.
We first perform a simulation study on a latent linear bandit with $d_{A}=50$ and $d_{K}=2$. We generate $\textbf{U}_{*}$ with $\textbf{U}_{ij}\text{ i.i.d. }\text{Unif}(0,\frac{2.5}{d_{K}d_{A}})$. We simulate the hidden labels $\bm{\theta}_{n}\sim\mathcal{N}(0,d_{K}^{-1}\textbf{I}_{d_{K}})$, generate feature vectors $\phi(x_{n,h},a_{n,h})\sim\mathcal{N}(0,\textbf{I}_{d_{A}})$ normalized to unit norm, and sample noise $\epsilon_{n,h}\text{ i.i.d. }\mathcal{N}(0,0.5^{2})$. We use SOLD to estimate $\hat{\textbf{U}}$ from the offline dataset $\mathcal{D}_{\operatorname{off}}$, which consists of $5000$ trajectories of length $20$ each. In accordance with the confidence set determined by Li et al. (2010), we choose $\alpha_{1,t}=0.33\sqrt{d_{K}\log(1+10T/d_{K})}$ and $\alpha_{2,t}=0.33\sqrt{d_{A}\log(1+10T/d_{A})}$, and share the LinUCB and ProBALL-UCB hyperparameters by assigning $\alpha_{t}=\alpha_{2,t}$.
Figure 2: Comparison of ProBALL-UCB with LinUCB, for different choices of $\tau$ and confidence bound constructions. All variants perform no worse than LinUCB, with martingale Bernstein performing the best. The shaded area depicts $1$-standard error confidence intervals over $30$ trials.
The results are presented in Figure 2. Note that ProBALL-UCB (Algorithm 3) performs no worse than LinUCB, no matter what we choose for $\tau$ and $\Delta_{\operatorname{off}}$. However, we see a distinct benefit from using tighter confidence bounds – as $\Delta_{\operatorname{off}}$ gets smaller, Algorithm 3 chooses to utilize the projected estimate $\hat{\bm{\beta}}_{1,t}$ more often, resulting in better performance. Note that the kinks in the regret curves correspond to points where Algorithm 3 switches over to the higher dimensional optimism in Line 9.
MovieLens dataset.
In line with Hong et al. (2020), we assess the performance of our algorithms on real data using the MovieLens dataset (Harper and Konstan, 2015). This is a large-scale movie recommendation dataset comprising $6040$ users and $3883$ movies, where each user may rate one or more movies. Like Hong et al. (2020), we filter the dataset to include only movies rated by at least $200$ users and vice-versa. We factor the sparse rating matrix into user parameters $\bm{\beta}$ and movie features $\Phi$ using the probabilistic matrix factorization algorithm of Mnih and Salakhutdinov (2007a), using nuclear norm regularization so that the rank of $\bm{\beta}$ is $d_{K}=18$. However, we consider a much higher dimensional problem than Hong et al. (2020) do – we let $d_{A}=200$ so $\bm{\beta}\in\mathbb{R}^{1589\times 200},\Phi\in\mathbb{R}^{200\times 1426}$. At each round for user $i$, the agent chooses between $20$ movies of different genres with features $\Phi_{a_{1}},...,\Phi_{a_{20}}$, and has to recommend the best movie presented to it to maximize the user’s rating of the movie. We generate rewards for recommending movie $j$ to user $i$ by $\bm{\beta}_{i}^{T}\Phi_{j}+\epsilon_{ij},\epsilon_{ij}\text{ i.i.d. }\mathcal{N}(0,0.5)$.
Our hyperparameters are chosen and varied just as in the simulation study. To reproduce the methods of Hong et al. (2020), we cluster the user features into $d_{K}$ clusters using k-means, and provide mUCB and mmUCB with the mean vectors of each cluster as latent models. We initialize ProBALL-UCB with a subspace estimated with an unregularized variant of SOLD, that uses pseudo-inverses instead of inverses, because of difficulties in finding an appropriate regularization parameter for this large, noisy, and high-dimensional dataset. The subspace was estimated from $5000$ trajectories of length $50$ simulated from the reward model and the uniform behavior policy. Note that we assign $\Delta_{\operatorname{off}}$ for $\epsilon$ in mmUCB, as this is their tolerance parameter for model misspecification.
Figure 3: Comparison of ProBALL-UCB initialized with SOLD against LinUCB, mUCB, and mmUCB, for different choices of $\tau$ and confidence bound constructions. All variants of ProBALL-UCB perform no worse than LinUCB, and outperform mUCB and mmUCB. Shaded area depicts $1$-standard error confidence intervals over $30$ trials with fresh $\bm{\theta}$. The confidence intervals on regret thus account for the variation in frequentist regret for changing $\bm{\theta}$.
Figure 3 depicts the result of this experiment. Once again, ProBALL-UCB performs no worse than LinUCB, no matter what we choose for $\tau$ and $\Delta_{\operatorname{off}}$, and the benefit of using tighter confidence bounds remains. With $\tau=0.5$, ProBALL-UCB with martingale Bernstein confidence bands stops using the projected estimates at around timestep $70$, but still continues to outperform LinUCB. Although mUCB and mmUCB perform slightly better than ProBALL-UCB and LinUCB at the beginning, the model misspecification incurred by discretizing the features into $d_{K}$ clusters ensures that it typically suffers linear regret in this scenario. The lower initial performance of ProBALL-UCB and Lin-UCB is a consequence of their higher initial exploration.
8 Discussion, Limitations and Further Work
In this paper, we have addressed the problem of leveraging offline data to accelerate online learning in linear latent bandits. Our work has a few limitations. First, while ProBALL-UCB is practical and computationally efficient, it has a slightly weaker worse-case guarantee than the less practical LOCAL-UCB. Second, when the size $N$ of the offline dataset is small, it can be hard to tune the regularization $\mu$. We use a pseudoinverse-based version of SOLD in such a case, implemented in our code. Third, the offline uncertainty sets computed using $\Delta_{\operatorname{off}}$ can be overly conservative, and a discount hyperparameter $\tau$ must be fine-tuned online.
Despite these limitations, our work enjoys strong theoretical guarantees and convincing empirical performance. We hope that this method opens the door for developing other efficient and scalable algorithms for sequential decision-making with continuous latent states. One can use ideas presented in this paper to design similar algorithms for MDPs, linear MDPs as well as under general function approximation for both MDPs and bandits. It is also possible that our regret bounds can be sharpened using feature-space partition ideas from hybrid RL (Li et al., 2023; Tan and Xu, 2024). Finally, it will be of interest to obtain minimax guarantees for leveraging offline data to speed up online learning in latent and linear latent contextual bandits.
Appendix A Additional Lemmas and Discussion
A.1 Broader Impacts
Often, protected groups with private identities have temporal behavior correlated with their private identities. Latent state estimation methods have the potential to identify such private identities online, and must be used with care. Use of such methods should comply with the relevant privacy and data protection acts, and corporations with access to a large customer base as well as governments should be mindful of the impact of the use of latent state methods on their customers and citizens respectively.
A.2 Coherence with Equality is Necessary
Lemma 2 (Coherence with Equality is Necessary).
There exists a stateless decision process that is exchangeable and not coherent but satisfies the following property: for any two action sequences $\tau,\tau^{\prime}$ of lengths $H$ and $H^{\prime}$ sharing the same actions $(a_{h},\dots a_{k})$ from index $h$ to $k$, with $\mathcal{F}_{H}(\tau)=(Y_{1},\dots Y_{H})$ and $\mathcal{F}_{H^{\prime}}(\tau^{\prime})=(Y_{1}^{\prime},\dots Y_{H^{\prime}}^{\prime})$, we have $(Y_{h},\dots Y_{k})\sim(Y_{h}^{\prime},\dots Y_{k}^{\prime})$. Additionally, since the SDP is not coherent, it is not a latent bandit.
Proof.
Consider an SDP with action set $\mathcal{A}=\{0,1\}$ equipped with a random variable $\theta\in\{0,1\}$ distributed as $Ber(1/2)$. If the first action is 00, then all rewards are given by $\theta$. If the first action is $1$, then the first reward $X_{1}=\theta$ while all future rewards are $1-\theta$.
Notice now that for any timestep $h$, its rewards are $Ber(1/2)$. We note that this process is clearly exchangeable. Moreover, this process is coherent in a weaker sense – for any two action sequences $\tau,\tau^{\prime}$ of lengths $H$ and $H^{\prime}$ sharing the same actions $(a_{h},\dots a_{k})$ from index $h$ to $k$, with $\mathcal{F}_{H}(\tau)=(Y_{1},\dots Y_{H})$ and $\mathcal{F}_{H^{\prime}}(\tau^{\prime})=(Y_{1}^{\prime},\dots Y_{H^{\prime}}^{\prime})$, we have $(Y_{h},\dots Y_{k})\sim(Y_{h}^{\prime},\dots Y_{k}^{\prime})$. That is, the reward subsequences are only equal in distribution. This is easy to see, since if $a_{1}$ is not included in the action subsequence or if $a_{1}=0$, the reward subsequence is always given by $(X,\dots X)$ for $X\sim Ber(1/2)$. If $a_{1}=1$ and $a_{1}$ is in the action subsequence, then the reward subsequence is always given by $(-X,X,\dots X)$ for $X\sim Ber(1/2)$.
However, the reward subsequences are not equal in value. If the action subsequence is inside two action sequences $\tau=(a_{1},\dots a_{H})$ and $\tau^{\prime}=(a^{\prime}_{1},\dots a^{\prime}_{H})$ with $a_{1}=0$ and $a_{1}^{\prime}=1$, then their rewards are negatives of each other. So, this SDP is not coherent. This means that this SDP is not a latent bandit. ∎
A.3 Contexts, Latent State and Behavior Policy cannot be Dependent
See 1
Proof.
Let $e_{1},e_{2}$ be the standard basis of $\mathbb{R}^{2}$. For all these examples, our two latent bandits will satisfy the following:
  • Latent contextual bandit $1$: Let $\bm{\theta}$ take values $e_{1}+e_{2}$ and $-e_{1}-e_{2}$ with probability $1/2$ each.
  • Latent contextual bandit $2$: Let $\bm{\theta}$ take values $e_{1}-e_{2}$ and $e_{2}-e_{1}$ with probability $1/2$ each.
A.3.1 Contexts cannot be dependent on each other
We consider two actions, so $\mathcal{A}=\{0,1\}$ We have two contexts $x,y$ so that $\phi(x,0)=e_{1}$, $\phi(x,1)=2e_{1}$ while $\phi(y,0)=e_{2}$, $\phi(y,1)=2e_{2}$. We design them to be dependent so that any trajectory either only sees $x$ or only sees $y$. However, $x$ and $y$ are both seen with probability $1/2$. Pick $\pi_{b}$ so that it takes each action with probability $1/2$. Now note that the mean reward of $x,0$ in either latent bandit takes values ±$\pm 1$ with probability $1/2$. So, the offline data distributions are also indistinguishable and every context-action pair is seen with probability at least $1/4$.
A.3.2 Contexts and latent state cannot be dependent
We consider two actions, so $\mathcal{A}=\{0,1\}$ Again, consider two contexts $x,y$ so that $\phi(x,0)=e_{1}$, $\phi(x,1)=2e_{1}$ while $\phi(y,0)=e_{2}$, $\phi(y,1)=2e_{2}$. Let us say that for either latent bandit and for any $\bm{\theta}$, the context distribution is a Dirac-$\delta$ over the context whose features have a positive dot product with $\bm{\theta}$. Then, note that either context is seen in both latent bandits with probability $1/2$. Again, let $\pi_{b}$ take either action with probability $1/2$. So, the offline data distributions are indistinguishable and every context-action pair is seen with probability at least $1/4$.
A.3.3 Latent state and behavior policy cannot be dependent
We consider four actions, so that $\mathcal{A}=\{0,1,2,3\}$. Let there be a single context $x$ with $\phi(x,0)=e_{1},\phi(x,1)=e_{2},\phi(x,2)=-e_{1},\phi(x,3)=-e_{2}$. For any latent state $\bm{\theta}$ in either bandit, let $\pi_{b}$ be uniform over the actions that have a positive dot product with $\bm{\theta}$. It is then easy to see that the offline data distributions are indistinguishable and every context-action pair is seen with probability at least $1/4$. ∎
Appendix B A de Finetti Theorem for decision processes
B.1 Discussion on Liu et al. [2023]’s smaller class of generalized bandits
We prove the de Finetti theorem for a very general formulation of decision processes. However, past work [Liu et al., 2023] has studied a simpler generalization of bandits, namely a stochastic process valued in $\mathbb{R}^{\mathcal{A}}$. A sample point in this space is a sequence of functions in $\mathcal{A}\to\mathbb{R}$, which rules out the possibility of adaptivity. In contrast, a sample point in our space is a function of sequences, which subsumes all sample points of their space, but allows for adaptivity.
Liu et al. [2023] are able to use the original de Finetti theorem on their random process directly, but work with a much more restrictive kind of decision process. We show that even when considering much more general stateless decision processes, latent bandits are the "right" objects produced by a de Finetti theorem for stateless decision processes.
B.2 Proof of the De Finetti Theorem for Stateless Decision Processes
A note on $F$, measurability and well-definedness in the definition of a latent bandit.
Recall that $F$ in the definition of a latent bandit needs to be a measurable map $\Omega\to(\mathcal{A}\to\mathcal{P}(\mathbb{R}))$. To define measurability for the output space of functions $(\mathcal{A}\to\mathcal{P}(\mathbb{R}))$, we endow $\mathcal{P}(\mathbb{R})$ with the topology of weak convergence, endow the space $\mathcal{P}(\mathbb{R})^{\mathcal{A}}$ of maps $\mathcal{A}\to\mathcal{P}(\mathbb{R})$ with the topology of point-wise convergence, and require $F$ to be measurable w.r.t. the induced Borel $\sigma$-algebra. We also recall two abuses of notation in the definition of a latent bandit. First, we abuse notation to define the random measure $F(a):=(\omega\mapsto F(\omega)(a))$. Second, we also abuse notation to conflate $F(a)$ with the curried map $\kappa_{a}(\omega,B):=F(a)(\omega)(B)$, which is a map $\Omega\times\mathcal{B}\to[0,1]$. This map $\kappa_{a}$ will turn out to be a kernel by the construction of $F$. Equating $F(a)$ to a regular conditional distribution in the definition of a latent bandit requires that $\kappa_{a}$ be a kernel.
We recall our de Finetti theorem for stateless decision processes here. See 1
Proof.
Consider an exchangeable and coherent stateless decision process. We will establish that there is a latent bandit with the same reward distribution as this process for any sequence of actions.
For any sequence $\tau=(a_{1},\dots a_{H})$, denote by $(Y_{\tau,1},\dots Y_{\tau,H}):=\mathcal{F}_{H}(a_{1},\dots a_{H})$. We intend to establish that there is a random measure-valued function $F$ such that $(Y_{\tau,1},\dots Y_{\tau,H})$ are independent given $F$ and $\mathcal{L}[Y_{\tau,h}\mid F]=F(a_{h})$ for all $h\leq H$ almost surely. Since conditional independence is a property of finite subsets of a set of random variables, it suffices to show this for finite $H$. The version for $H=\infty$ will immediately follow by coherence.
First recall that regular conditional distributions $\mathcal{L}[Y\mid F]$ are almost surely unique under if the $\sigma$-algebra of the output space of $Y$ is countably generated. Since the Borel $\sigma$-algebra on $\mathbb{R}$ is countably generated, this is true for our case. We also recall for the rusty reader that conditioning on a random variable is the same as conditioning on the induced $\sigma$-algebra in the domain.
B.3 Constructing $F$
Fix any finite trajectory $\tau=(a_{1},\dots a_{H})$ and index $h$.
The trick: Consider the infinite sequence $\tau_{\infty}:=(a_{h},a_{h},\dots)$. By coherence, $Y_{\tau,h}=Y_{\tau_{\infty},h}$. Now $\tau_{\infty}$ is a sequence where exchanging any finite set of rewards preserves the reward distribution, since all actions are identical. Since $\mathbb{R}$ is locally compact, we can apply the usual de Finetti representation theorem (Theorem 12.26 in Klenke [2008]) to conclude that:
  • The random measure $\Xi_{a_{h}}=\operatorname{wlim}_{n\to\infty}\frac{1}{n}\sum_{j=1}^{n}\delta_{Y_{\tau_{\infty},j}}$ is well defined. Here $\operatorname{wlim}$ is the weak limit of measures.
  • The regular conditional distribution $\mathcal{L}[Y_{\tau_{\infty},j}\mid\Xi_{a_{h}}]=\Xi_{a_{h}}$ for all $j$.
Since $Y_{\tau,h}=Y_{\tau_{\infty},h}$, we conclude that the conditional distribution $\mathcal{L}[Y_{\tau,h}\mid\Xi_{a_{h}}]=\mathcal{L}[Y_{\tau_{\infty},h}\mid\Xi_{a_{h}}]=\Xi_{a_{h}}$.
Constructing $F$: For every action, consider such a random measure $\Xi_{a_{h}}$ and define a random measure-valued function $F:\Omega\to(\mathcal{A}\to\mathcal{P}(\mathbb{R}))$ in the following manner: for any $\omega\in\Omega$, define $F(\omega)(a):=\Xi_{a_{h}}(\omega)$. We know that for any $a$, $\Xi_{a_{h}}$ is measurable w.r.t. the topology of weak convergence on $\mathcal{P}(\mathbb{R})$. It is now tedious but straightforward to verify that $F$ is measurable w.r.t. the Borel $\sigma$-algebra generated by the topology of pointwise convergence on $\mathcal{P}(\mathbb{R})^{\mathcal{A}}$.
B.4 Establishing that $\mathcal{L}[Y_{\tau,h}\mid F]=F(a_{h})$ almost surely for all $h$
Again, fix any finite trajectory $\tau=(a_{1},\dots a_{H})$ and index $h$. Recall that we abuse notation to denote by $F(a)$ the random-measure $\omega\mapsto F(\omega)(a)$. In particular, for any measurable set $B\subset\mathbb{R}$, we conflate $F(a)(\omega)(B)=F(\omega)(a)(B)=F(a)(\omega,B)$, where the last equality holds since $F$ is a regular conditional distribution. Note that $F(a_{h})=\Xi_{a_{h}}$ by the construction of $F$. Since $\Xi_{a_{h}}=F(a_{h})$, we have $\mathcal{L}[Y_{\tau,h}\mid F(a_{h})]=F(a_{h})$. Thus, it suffices to show that $\mathcal{L}[Y_{\tau,h}\mid F]=\mathcal{L}[Y_{\tau,h}\mid F(a_{h})]$.
Lemma 3.
$\mathcal{L}[Y_{\tau,h}\mid F]=\mathcal{L}[Y_{\tau,h}\mid F(a_{h})]$ almost surely.
Proof.
First note that $F(a_{h})$ is measurable w.r.t. $F$. For showing this, view the set of maps $\mathcal{P}(\mathbb{R})^{\mathcal{A}}$ as the product set $\prod_{a^{\prime}}\mathcal{P}(\mathbb{R})_{a^{\prime}}$. Now merely note that $F^{-1}(E\times\prod_{a^{\prime}\neq a}\mathcal{P}(\mathbb{R})_{a^{\prime}})=(F(a_{h}))^{-1}(E)$ for any measurable subset $E\subset\mathcal{P}(\mathbb{R})_{a}$. Hence, $F(a_{h})$ is measurable w.r.t. $F$.
Let $\mathcal{B}$ be the Borel $\sigma$-algebra on $\mathbb{R}$. Now recall that the regular conditional distribution $\mathcal{L}[X\mid G]$ for a real-valued random variable $X$ is the almost surely unique kernel $\kappa_{G,X}:\Omega\times\mathcal{B}\to\mathbb{R}$ such that:
  • $\omega\to\kappa_{G,X}(\omega,B)$ is $G$-measurable for any set $B\in\mathcal{B}$
  • $B\to\kappa_{G,X}(\omega,B)$ is a probability measure on $\mathbb{R}$ for any sample point $\omega\in\Omega$.
  • For any measurable set $B\subset\mathbb{R}$ and any $G$-measurable set $A$,
[…]
We will show our claim using the definition and a.s. uniqueness of the regular conditional distribution in our case. Consider any $F$-measurable set $A\subset\Omega$ and Borel-measurable set $B\subset\mathbb{R}$, $B\in\mathcal{B}$. Denote by $\kappa_{F}:=\mathcal{L}[Y_{\tau,h}\mid F]$ and by $\kappa_{a_{h}}:=\mathcal{L}[Y_{\tau,h}\mid F(a_{h})]=F(a)$. Note that by the coherence and exchangeability in section B.3,
[…]
for all $j$. Averaging all these equations and taking a limit, we get that
[…]
$=\lim_{n\to\infty}\frac{1}{n}\sum_{j=1}^{n}\mathbb{E}[\mathds{1}_{B}(Y_{\tau_{\infty},j})\mathds{1}_{A}]$
$=\mathbb{E}\left[\left(\lim_{n\to\infty}\frac{1}{n}\sum_{j=1}^{n}\mathds{1}_{B}(Y_{\tau_{\infty},j})\right)\mathds{1}_{A}\right]$
where the last equality holds by the dominated convergence theorem, if the limit exists. To establish that the limit exists and compute it, we apply the usual de Finetti theorem – specificaly point $(i)$ of remark 12.27 in Klenke [2008] with $f$ set to the identity map. This gives us that $\lim_{n\to\infty}\frac{1}{n}\sum_{j=1}^{n}\mathds{1}_{B}(Y_{\tau_{\infty},j})=\mathbb{E}[\mathds{1}_{B}(Y_{\tau_{\infty},h})\mid\Xi_{a_{h}}]=\Xi_{a_{h}}(B)$, where $\Xi_{a_{h}}(B)$ is the random variable $\omega\mapsto\Xi_{a_{h}}(\omega)(B)$. This in turn satisfies $\Xi_{a_{h}}(\omega)(B)=F(a_{h})(\omega)(B)=\kappa_{a_{h}}(\omega,B)$. This establishes that for any $F$-measurable set $A$ and Borel set $B$,
[…]
. In conclusion, $\kappa_{a_{h}}$ satisfies:
  • $F(a_{h})(B):=\omega\to\kappa_{a_{h}}(\omega,B)$ is $F$-measurable for any set $B\in\mathcal{B}$ since $F(a_{h})(B)$ is $F(a_{h})$ measurable by definition of $\kappa_{a_{h}}$ and $F(a_{h})$ is $F$-measurable from above.
  • $B\to\kappa_{a_{h}}(\omega,B)$ is a probability measure on $\mathbb{R}$ for any sample point $\omega\in\Omega$ by definition of $\kappa_{a_{h}}$
  • For any measurable set $B\subset\mathbb{R}$ and any $F$-measurable set $A$, by the argument above,
[…]
By the definition as well as a.s. uniqueness of regular conditional distributions in our case, this establishes that $\mathcal{L}[Y_{\tau,h}\mid F]=\kappa_{F}=\kappa_{a_{h}}=\mathbb{E}[\mathbb{E}[Y_{\tau,h}\mid F(a_{h})]$ almost surely.
B.5 Establishing conditional independence of rewards
This is the trickier to establish. It suffices to show that for any finite length $h$ trajectory $\tau=(a_{1},\dots a_{H})$ and any tuple $(f_{1},\dots f_{H})$ of bounded measurable functions, $\mathbb{E}\left[\prod_{h=1}^{H}f_{h}(Y_{\tau,h})\mid F\right]=\prod_{h=1}^{H}\mathbb{E}[f_{h}(Y_{\tau,h})\mid F]$. Equivalently, it suffices to show that for any bounded $F$-measurable real-valued random variable $U$, the following holds.
$\mathbb{E}\left[U\prod_{h=1}^{H}f_{h}(Y_{\tau,h})\right]=\mathbb{E}\left[U\prod_{h=1}^{H}\mathbb{E}[f_{h}(Y_{\tau,h})\mid F]\right]$
We will show this by induction on $H$. This clearly holds for $H=1$ by section B.3 above. Now assume that this holds for $H=l-1$.
Replacing the last term, $f_{l}(Y_{\tau,l})$, by an empirical average: Fix any trajectory $\tau=(a_{1},\dots a_{l})$. First consider the infinite trajectory $\tau^{\prime}=(a_{1},\dots a_{l-1},a_{l},a_{l},a_{l},\dots)$. This creates a random sequence of rewards $Y_{\tau^{\prime},1},Y_{\tau^{\prime},2},\dots$. Define $\tau_{j}^{\prime}$ by switching indices $l$ and $l+j$ in $\tau^{\prime}$ and considering the first $l$ actions. In particular, $\tau_{j}^{\prime}$ gives the sequence of rewards $(Y_{\tau^{\prime},1},\dots Y_{\tau^{\prime},l-1},Y_{\tau^{\prime},l+j})$.
First note that $\tau_{0}^{\prime}=\tau$. By coherence, $(Y_{\tau,1},\dots Y_{\tau,l})=(Y_{\tau^{\prime},1},\dots Y_{\tau^{\prime},l})$. By exchangeability and coherence, $(Y_{\tau^{\prime},1},\dots Y_{\tau^{\prime},l-1},Y_{\tau^{\prime},l})\sim(Y_{\tau^{\prime},1},\dots Y_{\tau^{\prime},l-1},Y_{\tau^{\prime},l+j})$. This means that for any $j\geq 0$,
$\mathbb{E}\left[U\prod_{h=1}^{l}f_{h}(Y_{\tau,h})\right]=\mathbb{E}\left[U\left(\prod_{h=1}^{l-1}f_{h}(Y_{\tau^{\prime},h})\right)f_{l}(Y_{\tau^{\prime},l+j})\right]$
We can then consider the average over all these equations for $j=0\to n-1$ and get that for all $n\geq 1$,
$\mathbb{E}\left[U\prod_{h=1}^{l}f_{h}(Y_{\tau,h})\right]=\mathbb{E}\left[U\left(\prod_{h=1}^{l-1}f_{h}(Y_{\tau^{\prime},h})\right)\left(\frac{1}{n}\sum_{j=0}^{n-1}f_{l}(Y_{\tau^{\prime},l+j})\right)\right]$
Taking limits and using the dominated convergence theorem, we get that
$\mathbb{E}\left[U\prod_{h=1}^{l}f_{h}(Y_{\tau,h})\right]=\mathbb{E}\left[U\left(\prod_{h=1}^{l-1}f_{h}(Y_{\tau^{\prime},h})\right)\left(\lim_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}f_{l}(Y_{\tau^{\prime},l+j})\right)\right]$ (1)
if the limit on the right side exists.
Showing that the empirical average is the conditional expectation: Again, consider a different infinite trajectory $\tau_{l}=(a_{l},a_{l},\dots)$. Again, by coherence, $Y_{\tau_{l},l+j}=Y_{\tau^{\prime},l+j}$ for all $j\geq 0$. From the usual de Finetti theorem, specifically point (i) of remark 12.27 in Klenke [2008], we have that $\frac{1}{m}\sum_{j=1}^{m}f_{l}(Y_{\tau_{l},j})\to\mathbb{E}[f_{l}(Y_{\tau_{l},l})\mid F]$. We can then observe that by general properties of convergence, the following holds.
$\mathbb{E}[f_{l}(Y_{\tau_{l},l})\mid F]$ $=\lim_{m\to\infty}\frac{1}{m}\sum_{j=1}^{m}f_{l}(Y_{\tau_{l},j})=\lim_{m\to\infty}\frac{1}{m-l}\sum_{j=l}^{m-1}f_{l}(Y_{\tau_{l},j})$
$=\lim_{m\to\infty}\frac{1}{m-l}\sum_{j=l}^{m-1}f_{l}(Y_{\tau^{\prime},j})=\lim_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}f_{l}(Y_{\tau^{\prime},l+j})$
We can combine this with equation 1 to get that
$\mathbb{E}\left[U\prod_{h=1}^{l}f_{h}(Y_{\tau,h})\right]=\mathbb{E}\left[U\left(\prod_{h=1}^{l-1}f_{h}(Y_{\tau^{\prime},h})\right)\mathbb{E}[f_{l}(Y_{\tau_{l},l})\mid F]\right]$
Since both $U$ and $\mathbb{E}[f_{l}(Y_{\tau_{l},l})\mid F]$ are now $F$ measurable, we can apply the induction hypothesis to $\tau^{\prime}$ truncated at $l-1$ and conclude that
$\mathbb{E}\left[U\prod_{h=1}^{l}f_{h}(Y_{\tau,h})\right]=\mathbb{E}\left[U\mathbb{E}[f_{l}(Y_{\tau_{l},l})\mid F]\prod_{h=1}^{l-1}f_{h}(Y_{\tau,h})\right]=\mathbb{E}\left[U\left(\prod_{h=1}^{l}\mathbb{E}[f_{h}(Y_{\tau^{\prime},h})\mid F]\right)\right]$
Thus, the induction step holds and the claim holds for all finite $H$. We discussed at the beginning of the section how this implies conditional independence for $H=\infty$ as well. ∎
B.6 A de Finetti theorem for TACDPs
Definition 7.
A TACDP is said to be coherent if for any $h\leq k\leq H,H^{\prime}\in\mathbb{N}\cup\{\infty\}$ and for any two context-action sequences $\tau,\tau^{\prime}$ of lengths $H$ and $H^{\prime}$ sharing the same context-action pairs $((x_{h},a_{h}),\dots(x_{k},a_{k}))$ from index $h$ to $k$, with $\mathcal{F}_{H}(\tau)=(Y_{1},\dots Y_{H})$ and $\mathcal{F}_{H^{\prime}}(\tau^{\prime})=(Y_{1}^{\prime},\dots Y_{H^{\prime}}^{\prime})$, we have $(Y_{h},\dots Y_{k})=(Y_{h}^{\prime},\dots Y_{k}^{\prime})$, viewed as functions of $\Omega$.
Definition 8.
A stateless decision process is said to be exchangeable if for any permutation $\pi:[H]\to[H]$ and $\mathcal{F}_{H}((x_{1},a_{1}),\dots(x_{H},a_{H}))=(Y_{1},\dots Y_{H})$, we have $\mathcal{F}_{h}((x_{\pi(1)},a_{\pi(1)}),\dots(x_{\pi(H)},a_{\pi(H)}))\sim(Y_{\pi(1)},\dots Y_{\pi(H)})$.
Definition 9.
A latent contextual bandit is a stateless decision process equipped with a random measure-valued function $F:\Omega\to((\mathcal{X}\times\mathcal{A})\to\mathcal{P}(\mathbb{R}))$ so that for any $H$ and context-action sequence $((x_{1},a_{1}),\dots(x_{H},a_{H}))$, the rewards $(Y_{1},\dots Y_{H}):=\mathcal{F}_{H}((x_{1},a_{1}),\dots(x_{H},a_{H}))$ are independent conditioned on $F$. Moreover, the conditional distribution $\mathcal{L}[Y_{h}\mid F]=F((x_{h},a_{h}))$ for all $h\leq H$ [^6] We abuse notation twice here. First, we write $F((x_{h},a_{h})):=(\omega\mapsto F(\omega)((x_{h},a_{h})))$. Second, as the regular conditional distribution $\mathcal{L}[Y_{h}\mid F]$ is a kernel that maps from $\Omega\times\mathcal{B}\to\mathbb{R}$, we view $F((x_{h},a_{h}))$ as its curried map $(\omega,B)\mapsto F((x_{h},a_{h}))(\omega)(B)$. A discussion of issues like measurability and well-definedness is in Appendix B.2.
Theorem 5 (De Finetti Theorem for Stateless Decision Processes).
Every exchangeable and coherent TACDP is a latent contextual bandit.
Proof.
The proof is verbatim the same as that for Theorem 1 after merely replacing $\mathcal{A}$ with $\mathcal{X}\times\mathcal{A}$ and $a$ with $(x,a)$. ∎
Appendix C Proofs for SOLD
Recall that $\mu_{\theta}:=\mathbb{E}[\bm{\theta}]$ and define $\mu_{\beta}:=\mathbb{E}[\bm{\beta}]=\textbf{U}_{\star}\mu_{\theta}$. Also recall that $\Lambda:=\mathbb{E}[\bm{\theta}_{n}\bm{\theta}_{n}^{\top}]$. For a trajectory with index $n$, denote by $\bm{\beta}_{n}:=\textbf{U}_{\star}\bm{\theta}_{n}$. Denote by $X_{n}:=[\phi(x_{1},a_{1}),\dots\phi(x_{H},a_{H})]^{\top}$. Denote by $\eta_{n}:=[\epsilon_{n,1},\epsilon_{n,2}\dots\epsilon_{n,H}]^{\top}$ the vector of subgaussian noises with subgaussian parameter $\sigma^{2}$. Since rewards are bounded by $R$, we know that $\sigma^{2}\leq R^{2}$. Denote by $X_{n,i}$ and $\eta_{n,i}$ the action matrix and noise vector corresponding to the trajectory halves $\tau_{n,i}$ for $i=1,2$. Since rewards are bounded by $R$, $\epsilon_{n,h}$ is subgaussian for all $h$ and so $\eta_{n,i}$ is $\sigma^{2}$-subgaussian for $i=1,2$. Also note that
$\hat{\bm{\beta}}_{n,i}$ $=(\mu I+X_{n,i}^{\top}X_{n,i})^{-1}X_{n,i}^{\top}r_{n,i}$
$=(\mu I+X_{n,i}^{\top}X_{n,i})^{-1}X_{n,i}^{\top}(X_{n,i}\bm{\beta}_{n}+\eta_{n,i})$
$=(I-\mu(\mu I+X_{n,i}^{\top}X_{n,i})^{-1})\bm{\beta}_{n}+(\mu I+X_{n,i}^{\top}X_{n,i})^{-1}X_{n,i}^{\top}\eta_{n,i}$
$=(I-\mu(\mu I+X_{n,i}^{\top}X_{n,i})^{-1})\bm{\beta}_{n}+(\mu I+X_{n,i}^{\top}X_{n,i})^{-1}X_{n,i}^{\top}\eta_{n,i}$
Note that $\hat{\bm{\beta}}_{n,1}$ and $\hat{\bm{\beta}}_{n,2}$ are identically distributed. Now recall that
$\textbf{M}_{n}=\frac{1}{2}(\hat{\bm{\beta}}_{n,1}\hat{\bm{\beta}}_{n,2}^{\top}+\hat{\bm{\beta}}_{n,2}\hat{\bm{\beta}}_{n,1}^{\top})$
are i.i.d. random matrices. Denote by $\textbf{D}_{n,i}:=(I-\mu(\mu I+X_{n,i}^{\top}X_{n,i})^{-1})$ and denote by $u_{i}:=(\mu I+X_{n,i}^{\top}X_{n,i})^{-1}X_{n,i}^{\top}\eta_{n,i}$.
Lemma 4.
If the per-reward noise is $\sigma^{2}$-subgaussian, then the following inequalities hold
$\|\textbf{M}_{n}\|_{2}$ $\leq R^{2}\left(2+\frac{H}{2\mu}\right)$
$\|\mathbb{E}[\textbf{M}_{n}^{2}]\|_{2}$ $\leq R^{4}+\frac{\sigma^{4}(d_{A}+1)}{8\mu^{2}}$
Proof.
We prove various bounds and assemble them.
Bounding $\|(\mu I+X_{n,i}^{\top}X_{n,i})^{-1}X_{n,i}^{\top}\|_{2}$: Consider any $X$ with SVD $X=U^{\top}\Sigma V$. This means that
$\|(\mu I+X^{\top}X)^{-1}X^{\top}\|_{2}$ $=V^{\top}(\Sigma^{2}+\mu)^{-1}\Sigma U$
$=\|V^{\top}(\Sigma^{2}+\mu)^{-1}\Sigma\|_{2}=\|(\Sigma^{2}+\mu)^{-1}\Sigma\|_{2}$
$\leq\max_{a}\frac{a}{a^{2}+\mu}$
$=\frac{1}{2\sqrt{\mu}}$
We can now apply this to $X=X_{n,i}$ and conclude that $\|(\mu I+X_{n,i}^{\top}X_{n,i})^{-1}X_{n,i}^{\top}\|_{2}\leq\frac{1}{2\sqrt{\mu}}$
$u_{i}$ are independent, $\frac{\sigma^{2}}{4\mu}$-subgaussian and $\frac{\sigma\sqrt{H}}{2\sqrt{\mu}}$-bounded: We claim that $u_{i}$ are independent, $\frac{\sigma^{2}}{4\mu}$-subgaussian and $\|u_{i}\|_{2}^{2}\leq\frac{\sigma\sqrt{H}}{2\sqrt{\mu}}$. Recall that $\eta_{n,i}$ are $\sigma^{2}$-subgaussian vectors and $\|(\mu I+X_{n,i}^{\top}X_{n,i})^{-1}X_{n,i}^{\top}\|_{2}\leq\frac{1}{2\sqrt{\mu}}$. So, we have that $u_{i}=(\mu I+X_{n,i}^{\top}X_{n,i})^{-1}X_{n,i}^{\top}\eta_{n,i}$ is $\frac{\sigma^{2}}{4\mu}$-subgaussian. Also recall that $u_{1}$ and $u_{2}$ are independent since both contexts and reward-noise are generated independently at each timestep. Finally, since $|\epsilon_{n,h}|\leq R$, we also have that $\|\eta_{n,i}\|_{2}^{2}\leq R^{2}H$, so $\|u_{i}\|_{2}^{2}\leq\frac{R^{2}H}{4\mu}$ since $\|(\mu I+X_{n,i}^{\top}X_{n,i})^{-1}X_{n,i}^{\top}\|_{2}\leq\frac{1}{2\sqrt{\mu}}$.
Bounding $\|\textbf{M}_{n}\|_{2}$: Note that $\|\textbf{D}_{n,i}\|_{2}\leq 1$, $\|\bm{\beta}_{2}\|_{n}\leq R$ and $\|u_{i}\|_{2}\leq\frac{\sigma\sqrt{H}}{2\sqrt{\mu}}$. So, we have that
$\|\textbf{M}_{n}\|_{2}\leq\left(R+\frac{R\sqrt{H}}{2\sqrt{\mu}}\right)^{2}\leq R^{2}\left(2+\frac{H}{2\mu}\right)$
Bounding $\|\mathbb{E}[\textbf{M}_{n}^{2}]\|_{2}$: We compute that
$\mathbb{E}[\textbf{M}_{n}^{2}]$ $=\frac{1}{4}\Big{(}\mathbb{E}[\textbf{D}_{n,1}\bm{\beta}_{n}\bm{\beta}_{n}^{\top}\textbf{D}_{n,1}]\mathbb{E}[\bm{\beta}_{n}^{\top}\textbf{D}_{n,2}^{2}\bm{\beta}_{n}]+\mathbb{E}[\textbf{D}_{n,2}\bm{\beta}_{n}\bm{\beta}_{n}^{\top}\textbf{D}_{n,2}]\mathbb{E}[\bm{\beta}_{n}^{\top}\textbf{D}_{n,1}^{2}\bm{\beta}_{n}]$
$\qquad+\mathbb{E}[(\bm{\beta}_{n}^{\top}\textbf{D}_{n,2}\textbf{D}_{n,1}\bm{\beta}_{n})\textbf{D}_{n,1}\bm{\beta}_{n}\bm{\beta}_{n}^{\top}\textbf{D}_{n,2}]]+\mathbb{E}[(\bm{\beta}_{n}^{\top}\textbf{D}_{n,1}\textbf{D}_{n,2}\bm{\beta}_{n})\textbf{D}_{n,2}\bm{\beta}_{n}\bm{\beta}_{n}^{\top}\textbf{D}_{n,1}]]$
$\qquad\mathbb{E}[\|u_{1}\|_{2}^{2}]\mathbb{E}[u_{2}u_{2}^{\top}]+\mathbb{E}[\|u_{2}\|_{2}^{2}]\mathbb{E}[u_{1}u_{1}^{\top}]+\mathbb{E}[u_{1}u_{2}^{\top}u_{1}u_{2}^{\top}]+\mathbb{E}[u_{2}u_{1}^{\top}u_{2}u_{1}^{\top}]\Big{)}$
Now since $\|\textbf{D}_{n,i}\|_{2}\leq 1$ and $\|\beta_{n}\|_{2}\leq R$, the norm of the first four terms is bounded by $R^{4}$. Now note that
$\|\mathbb{E}[u_{i}u_{i}^{\top}]\|_{2}=\max_{\textbf{v},\|\textbf{v}\|_{2}\leq 1}\mathbb{E}[\textbf{v}^{\top}u_{i}u_{i}^{\top}\textbf{v}]=\max_{\textbf{v},\|\textbf{v}\|_{2}\leq 1}\mathbb{E}[(u_{i}^{\top}\textbf{v})^{2}]\leq\frac{\sigma^{2}}{2\mu}$
$\|\mathbb{E}[\|u_{i}\|_{2}^{2}]\|_{2}=\operatorname{Tr}(\mathbb{E}[u_{i}u_{i}^{\top}])\leq\|\mathbb{E}[u_{i}u_{i}^{\top}]\|_{2}d_{A}\leq\frac{\sigma^{2}d_{A}}{2\mu}$
$\|\mathbb{E}[u_{2}u_{1}^{\top}u_{2}u_{1}^{\top}]\|_{2}=\|\mathbb{E}[u_{1}u_{2}^{\top}u_{1}u_{2}^{\top}]\|_{2}=\max_{\textbf{v},\|\textbf{v}\|_{2}\leq 1}\mathbb{E}[\textbf{v}^{\top}u_{1}u_{2}^{\top}u_{1}u_{2}^{\top}\textbf{v}]$
$\max_{\textbf{v},\|\textbf{v}\|_{2}\leq 1}\mathbb{E}[\textbf{v}^{\top}u_{1}u_{2}^{\top}u_{1}u_{2}^{\top}\textbf{v}]=\max_{\textbf{v},\|\textbf{v}\|_{2}\leq 1}\operatorname{Tr}(\textbf{v}^{\top}\mathbb{E}[u_{1}u_{1}^{\top}]\mathbb{E}[u_{2}u_{2}^{\top}]\textbf{v}]=\|\mathbb{E}[u_{1}u_{1}^{\top}]\mathbb{E}[u_{2}u_{2}^{\top}]\|_{2}\leq\frac{\sigma^{4}}{4\mu^{2}}$
Combining all of these, we get that
$\|\mathbb{E}[\textbf{M}_{n}^{2}]\|_{2}\leq R^{4}+\frac{\sigma^{4}(d_{A}+1)}{8\mu^{2}}$
Proposition 1 (Confidence Bound for $\overline{\textbf{M}}_{N}$).
With probability at least $21-\delta/$ 2, we have that $\|\overline{\textbf{M}}_{N}-\mathbb{E}[\textbf{M}_{1}]\|_{2}\leq\Delta_{M}$ with
$\Delta_{M}$ $:=\sqrt{2\left\|\sum_{n=1}^{N}\textbf{M}_{n}^{2}\right\|_{2}\frac{\log(4d_{A}/\delta)}{N}}+2R^{2}\left(2+\frac{H}{2\mu}\right)\left(\frac{2\log(4d_{A}/\delta)}{N}\right)^{3/4}$
$\qquad+4R^{2}\left(2+\frac{H}{2\mu}\right)\frac{\log(4d_{A}/\delta)}{3N}$
Proof.
Now since $\|\textbf{M}_{n}^{2}\|_{2}\leq\|\textbf{M}_{n}\|_{2}^{2}\leq R^{4}\left(2+\frac{H}{4\mu}\right)^{2}$, we have that $\|\textbf{M}_{n}^{2}-\mathbb{E}[\textbf{M}_{1}^{2}]\|_{2}\leq 2R^{4}\left(2+\frac{H}{4\mu}\right)^{2}$ by the matrix Hoeffding bound, we have that with probability $21-\delta/$ 2,
$\left\|\frac{1}{N}\sum_{n=1}^{N}\textbf{M}_{n}^{2}-\mathbb{E}[\textbf{M}_{1}^{2}]\right\|_{2}\leq 4R^{4}\left(2+\frac{H}{2\mu}\right)^{2}\sqrt{\frac{\log(d_{A}/\delta)}{N}}$
Further, by the matrix Bernstein inequality, we have that with probability $21-\delta/$ 2,
$\left\|\frac{1}{N}\sum_{n=1}^{N}\textbf{M}_{n}-\mathbb{E}[\textbf{M}_{1}]\right\|_{2}\leq\sqrt{2\|\mathbb{E}[\textbf{M}_{1}^{2}]\|_{2}\frac{\log(d_{A}/\delta)}{N}}+4R^{2}\left(2+\frac{H}{2\mu}\right)\frac{\log(d_{A}/\delta)}{3N}$
Combining the two results and using a union bound, we get that with probability $21-\delta/$ 2
$\|\overline{\textbf{M}}_{N}-\mathbb{E}[\textbf{M}_{1}]\|_{2}$ $=\left\|\frac{1}{N}\sum_{n=1}^{N}\textbf{M}_{n}-\mathbb{E}[\textbf{M}_{1}]\right\|_{2}$
$\leq\sqrt{2\left\|\sum_{n=1}^{N}\textbf{M}_{n}^{2}\right\|_{2}\frac{\log(2d_{A}/\delta)}{N}}+2R^{2}\left(2+\frac{H}{2\mu}\right)\left(\frac{2\log(2d_{A}/\delta)}{N}\right)^{3/4}$
$\qquad+4R^{2}\left(2+\frac{H}{2\mu}\right)\frac{\log(d_{A}/\delta)}{3N}$
Proposition 2 (Confidence Bound for $\overline{\textbf{D}}_{N,i}$).
With probability $41-\delta/$ 4, for $i=1,2$, we have that $\|\overline{\textbf{D}}_{N,i}-\mathbb{E}[\textbf{D}_{n,i}]\|_{2}=\left\|\frac{1}{N}\sum_{n=1}^{N}\textbf{D}_{n,i}-\mathbb{E}[\textbf{D}_{n,i}]\right\|_{2}\leq\Delta_{D}$ with
$\Delta_{D}\leq\sqrt{\frac{8\log(4d_{A}/\delta)}{N}}$
Proof.
Since $\|\textbf{D}_{n,i}\|_{2}\leq 1$, this immediately follows by the matrix Hoeffding inequality. ∎
Lemma 5 (Confidence Bound for $\overline{\textbf{D}}_{N,1}^{-1}\overline{\textbf{M}}_{N}\overline{\textbf{D}}_{N,2}^{-1}$).
We have that with probability $1-\delta$
$\|\overline{\textbf{D}}_{N,1}^{-1}\overline{\textbf{M}}_{N}\overline{\textbf{D}}_{N,2}^{-1}-\mathbb{E}[\textbf{D}_{N,1}]^{-1}\mathbb{E}[\textbf{M}_{N,1}]\mathbb{E}[\textbf{D}_{N,2}]^{-1}\|_{2}$ $\leq\left(\frac{B_{D}^{3}(2-B_{D}\Delta_{D})}{(1-B_{D}\Delta_{D})^{2}}\right)(R^{2}+\Delta_{M})\Delta_{D}$
$\qquad+\left(\frac{B_{D}}{1-B_{D}\Delta_{D}}\right)^{2}\Delta_{M}$
where $B_{D}=\max_{i=1,2}\|\overline{\textbf{D}}_{N,i}^{-1}\|_{2}$.
Proof.
For brevity, just for this proof, we define $\textbf{D}_{i}=\mathbb{E}[\overline{\textbf{D}}_{N,i}]$ for $i=1,2$ and by $\textbf{M}:=\mathbb{E}[\textbf{M}_{1}]=\mathbb{E}[\overline{\textbf{M}}_{N}]$. By a union bound, Propositions 1 and 2 hold with probability at least $1-\delta$. The statements in the rest of this proof thus hold with probability at least $1-\delta$. Now note that
$\|\overline{\textbf{D}}_{N,1}^{-1}\overline{\textbf{M}}_{N,1}\overline{\textbf{D}}_{N,2}^{-1}-\textbf{D}_{1}^{-1}\textbf{M}\textbf{D}_{2}^{-1}\|_{2}$ $\leq\|\overline{\textbf{D}}_{N,1}^{-1}\|_{2}\|\overline{\textbf{M}}_{N,1}\|_{2}\|\overline{\textbf{D}}_{N,2}^{-1}-\textbf{D}_{2}^{-1}\|_{2}$
$\qquad+\|\overline{\textbf{D}}_{N,1}^{-1}-\textbf{D}_{1}^{-1}\|_{2}\|\overline{\textbf{M}}_{N,1}\|_{2}\|\textbf{D}^{-1}\|_{2}$
$\qquad+\|\textbf{D}_{1}^{-1}\|_{2}\|\overline{\textbf{M}}_{N,1}-\textbf{M}\|_{2}\|\textbf{D}_{2}^{-1}\|_{2}$
Now note that by inequality (1.1) from Wei et al. [2005], we have that
$\|\overline{\textbf{D}}_{N,i}^{-1}-\textbf{D}_{i}^{-1}\|_{2}\leq\frac{B_{D}^{2}\Delta_{D}}{1-B_{D}\Delta_{D}}$
This means that
$\|\textbf{D}_{i}^{-1}\|_{2}\leq\|\overline{\textbf{D}}_{N,i}^{-1}\|_{2}+\|\overline{\textbf{D}}_{N,i}^{-1}-\textbf{D}_{i}^{-1}\|_{2}\leq\frac{B_{D}}{1-B_{D}\Delta_{D}}$
Also, since contexts in both trajectory halves have the same distribution and contexts are generated independently, $\textbf{M}=\mathbb{E}[\overline{\textbf{M}}_{1}]=\textbf{D}_{i}\mathbb{E}[\beta_{1}\beta_{1}^{\top}]\textbf{D}_{i}$. So, $\|\textbf{M}\|_{2}\leq\|\textbf{D}_{1}\|_{2}\|\textbf{D}_{2}\|_{2}R^{2}\leq R^{2}$. This implies that
$\|\overline{\textbf{M}}_{N,1}\|_{2}\leq R^{2}+\|\overline{\textbf{M}}_{N,1}-\textbf{M}\|_{2}\leq R^{2}+\Delta_{M}$
Combining all these with the bound above, we get that
$\|\overline{\textbf{D}}_{N,1}^{-1}\overline{\textbf{M}}_{N}\overline{\textbf{D}}_{N,2}^{-1}-\mathbb{E}[\overline{\textbf{D}}_{N,1}]^{-1}\mathbb{E}[\textbf{M}_{n,1}]\mathbb{E}[\overline{\textbf{D}}_{N,2}]^{-1}\|_{2}$ $\leq\left(\frac{B_{D}^{3}(2-B_{D}\Delta_{D})}{(1-B_{D}\Delta_{D})^{2}}\right)(R^{2}+\Delta_{M})\Delta_{D}$
$\qquad+\left(\frac{B_{D}}{1-B_{D}\Delta_{D}}\right)^{2}\Delta_{M}$
See 2
Proof.
First note that $\mathbb{E}[\textbf{D}_{N,1}]^{-1}\mathbb{E}[\textbf{M}_{N,1}]\mathbb{E}[\textbf{D}_{N,2}]^{-1}=\textbf{U}_{\star}(\Lambda+\mu_{\theta}\mu_{\theta}^{\top})\textbf{U}_{\star}^{\top}$. Also recall that $\hat{\textbf{U}}$ is given by the top-$d_{K}$ eigenvectors for $\mathbb{E}[\textbf{D}_{N,1}^{-1}\textbf{M}_{N,1}\textbf{D}_{N,2}^{-1}]$. This means that there is a $d_{K}\times d_{K}$ unitary matrix W such that $\textbf{U}_{\star}\textbf{W}$ forms the eigenvectors for $\textbf{U}_{\star}(\Lambda+\mu_{\theta}\mu_{\theta}^{\top})\textbf{U}_{\star}^{\top}$. This means that by the Davis-Kahan theorem for statisticians [Yu et al., 2014] and Lemma 5, we have that with probability $1-\delta$
$\|\hat{\textbf{U}}\hat{\textbf{U}}^{\top}-\textbf{U}_{\star}\textbf{U}_{\star}^{\top}\|_{2}$ $=\|\hat{\textbf{U}}\hat{\textbf{U}}^{\top}-\textbf{U}_{\star}\textbf{W}\textbf{W}^{\top}\textbf{U}_{\star}^{\top}\|_{2}=\sqrt{2}(d_{K}-\|\hat{\textbf{U}}^{\top}\textbf{U}_{\star}\|_{F}^{2})$
$\leq\frac{2\sqrt{2d_{K}}}{\hat{\lambda}}\|\mathbb{E}[\textbf{D}_{N,1}]^{-1}\mathbb{E}[\textbf{M}_{N,1}]\mathbb{E}[\textbf{D}_{N,2}]^{-1}-\textbf{D}_{N,1}^{-1}\textbf{M}_{N,1}\textbf{D}_{N,2}^{-1}\|_{2}$
$\leq\frac{2\sqrt{2d_{K}}}{\hat{\lambda}}\left(\left(\frac{B_{D}^{3}(2-B_{D}\Delta_{D})}{(1-B_{D}\Delta_{D})^{2}}\right)(R^{2}+\Delta_{M})\Delta_{D}+\left(\frac{B_{D}}{1-B_{D}\Delta_{D}}\right)^{2}\Delta_{M}\right)$
We thus set $\Delta_{\operatorname{off}}$ to be the value above.
Bounding $\Delta_{\operatorname{off}}$: Now note that for large enough $N$, $\|\mathbb{E}[\textbf{D}_{n,1}]^{-1}\|_{2}\Delta_{D}=\sqrt{\frac{8\log(d_{A}/\delta)}{\lambda_{A}N}}\leq\frac{1}{2}$. In particular, by applying inequality (1.1) from Wei et al. [2005], we have that $B_{D}\leq 2\|\mathbb{E}[\textbf{D}_{n,1}]^{-1}\|_{2}=\frac{2}{\lambda_{A}}$. This already gives us the much simpler expression
$\Delta_{\operatorname{off}}\leq\frac{2\sqrt{2d_{K}}}{\hat{\lambda}}\left(\frac{64}{\lambda_{A}^{3}}(R^{2}+\Delta_{M})\Delta_{D}+\frac{16}{\lambda_{A}^{2}}\Delta_{M}\right)$
Also note that by Lemma 4 and the matrix Hoeffding inequality, we have that
$\|\frac{1}{N}\sum_{n=1}^{N}\textbf{M}_{n}^{2}\|$ $\leq\|\mathbb{E}[\textbf{M}_{1}^{2}]\|+4R^{4}\left(2+\frac{H}{2\mu}\right)^{2}\sqrt{\frac{\log(d_{A}/\delta)}{N}}$
$\leq R^{4}+\frac{\sigma^{4}(d_{A}+1)}{8\mu^{2}}+4R^{4}\left(2+\frac{H}{2\mu}\right)^{2}\sqrt{\frac{\log(d_{A}/\delta)}{N}}$
We combine this with Proposition 1 to get that
$\Delta_{M}=O\left(\left(R^{2}+\frac{\sigma^{2}\sqrt{d_{A}}}{\mu}\right)\sqrt{\frac{\log(d_{A}/\delta)}{N}}\right)$
Since $\sigma^{2}\leq R^{2}$, we get that
$\Delta_{M}=O\left(R^{2}\sqrt{\frac{d_{A}\log(d_{A}/\delta)}{N}}\right)=O(R^{2})$
Also recall from Proposition 2, we get that
$\Delta_{D}=\sqrt{\frac{8\log(d_{A}/\delta)}{N}}$
Also recall that $\lambda_{\theta}$ is the minimum eigenvalue of $\frac{1}{R^{2}}\Lambda$, and so the minimum eigenvalue of $\textbf{U}_{\star}(\Lambda+\mu_{\theta}\mu_{\theta}^{\top})\textbf{U}_{\star}^{\top}$ is larger than $R^{2}\lambda_{\theta}$. We then conclude that $\hat{\lambda}\geq R^{2}\lambda_{\theta}-2\Delta_{M}\geq\frac{\lambda_{\theta}}{2}$ for large enough $N$. Combining all these, we get that
$\Delta_{\operatorname{off}}$ $=O\left(\frac{\sqrt{d_{K}}}{R^{2}\lambda_{\theta}}\left(\frac{R^{2}}{\lambda_{A}^{3}}+\frac{R^{2}}{\lambda_{A}^{2}}\sqrt{d_{A}}\right)\sqrt{\frac{\log(d_{A}/\delta)}{N}}\right)$
$=O\left(\frac{1}{\lambda_{\theta}\lambda_{A}^{3}}\sqrt{\frac{d_{K}d_{A}\log(d_{A}/\delta)}{N}}\right)$
Appendix D Proofs for LOCAL-UCB
See 3
Proof.
Let the true latent state for the given trajectory be $\bm{\theta}_{\star}$, so that the reward parameter $\bm{\beta}_{\star}=\textbf{U}_{\star}\bm{\theta}_{\star}$.
Showing that $\textbf{U}_{\star},\bm{\beta}_{\star}$ are in our confidence set: We check all our constraints:
  • Note that with probability $31-\delta/$ 3, $\textbf{U}_{\star}$ satisfies $\|\hat{\textbf{U}}^{\top}\textbf{U}_{\star}\|_{F}\geq\sqrt{d_{K}-\Delta_{\operatorname{off}}^{2}/2}$.
  • From the standard confidence ellipsoid bound for linear models applied to dimension $d_{A}$, with probability $1-\delta/3T$, $\|\hat{\bm{\beta}}_{2,t}-\bm{\beta}_{\star}\|_{\textbf{V}_{t}^{-1}}\leq\alpha_{2,t}$ for all $t$.
  • Note that $\textbf{U}_{\star}\textbf{U}_{\star}^{\top}\bm{\beta}_{\star}=\bm{\beta}_{\star}$.
  • Finally, we apply the standard confidence ellipsoid bound for linear models to dimension $d_{K}$ instead of $d_{A}$ with model $r_{t}=(\phi(x_{t},a_{t})^{\top}\textbf{U}_{\star})(\textbf{U}_{\star}^{\top}\beta_{\star})+\epsilon_{t}$. This means that $\|\textbf{U}_{\star}^{\top}\beta_{\star}-\textbf{U}_{\star}^{\top}\hat{\beta}_{1,t}\|_{\textbf{V}_{1,t}^{-1}}\leq\alpha_{1,t}$ where $\textbf{V}_{1,t}=(\textbf{I}_{d_{K}}+\sum_{s=1}^{t-1}\textbf{U}_{\star}^{\top}\phi(x_{t},a_{t})\phi(x_{t},a_{t})^{\top}\textbf{U}_{\star})$. Note that since $\textbf{U}_{\star}\textbf{U}_{\star}^{\top}=\textbf{I}_{d_{K}}$, we have that $\textbf{V}_{1,t}=(\textbf{U}_{\star}^{\top}\textbf{V}_{t}\textbf{U}_{\star})$. So, $\|\textbf{U}_{\star}^{\top}(\bm{\beta}-\hat{\bm{\beta}}_{1,t})\|_{(\textbf{U}_{\star}^{T}\textbf{V}_{t}\textbf{U}_{\star})^{-1}}\leq\alpha_{1,t}$ holds with probability at least $1-\delta/3T$ for all $t$.
So, by a union bound over all events, we get that for all actions $a$, the tuple $(a,\bm{\beta}_{\star},\textbf{U}_{\star})$ satisfies our conditions with probability $1-\delta$.
Leveraging optimism in low dimension: From above and by the optimistic design of the algorithm, with probability $1-\delta$, $\phi(x_{t},a_{t})^{\top}\tilde{\bm{\beta}}_{t}$ is an upper bound on $\phi(x_{t},a)^{\top}\bm{\beta}_{\star}$ for any action $a$. Thus, we have the following regret decomposition with probability at least $1-\delta$:
$\operatorname{Reg}_{T}$ $=\sum_{t=1}^{T}\phi(x_{t},a_{t}^{\star})^{\top}\bm{\beta}_{\star}-\phi(x_{t},a_{t})^{\top}\bm{\beta}_{\star}$
$\leq\sum_{t=1}^{T}\phi(x_{t},a_{t})^{\top}\tilde{\bm{\beta}}_{t}-\phi(x_{t},a_{t})^{\top}\bm{\beta}_{\star}$
=$\stackrel{{\scriptstyle(i)}}{{=}}\sum_{t=1}^{T}\phi(x_{t},a_{t})^{\top}\tilde{\textbf{U}}_{t}\tilde{\textbf{U}}_{t}^{\top}\tilde{\bm{\beta}}_{t}-\phi(x_{t},a_{t})^{\top}\textbf{U}_{\star}\textbf{U}_{\star}^{\top}\bm{\beta}_{\star}$
$\leq\sum_{t=1}^{T}\phi(x_{t},a_{t})^{\top}\tilde{(}\textbf{U}_{t}\tilde{\textbf{U}}_{t}^{\top}-\textbf{U}_{\star}\textbf{U}_{\star}^{\top})\tilde{\bm{\beta}}_{t}+\phi(x_{t},a_{t})^{\top}\textbf{U}_{\star}\textbf{U}_{\star}^{\top}(\bm{\beta}_{t}-\bm{\beta}_{\star})$
$\leq RT\|\textbf{U}_{t}\tilde{\textbf{U}}_{t}^{\top}-\textbf{U}_{\star}\textbf{U}_{\star}^{\top}\|_{2}+\sum_{t=1}^{T}\phi(x_{t},a_{t})^{\top}\textbf{U}_{\star}\textbf{U}_{\star}^{\top}(\bm{\beta}_{t}-\bm{\beta}_{\star})$
$\leq RT\Delta_{\operatorname{off}}+\sum_{t=1}^{T}\phi(x_{t},a_{t})^{\top}\textbf{U}_{\star}\textbf{U}_{\star}^{\top}(\bm{\beta}_{t}-\bm{\beta}_{\star})$
$\leq RT\Delta_{\operatorname{off}}+\sum_{t=1}^{T}\|\phi(x_{t},a_{t})^{\top}\textbf{U}_{\star}\|_{(\textbf{U}_{\star}^{\top}\textbf{V}_{t}\textbf{U}_{\star})^{-1}}\|\textbf{U}_{\star}^{\top}(\bm{\beta}_{t}-\bm{\beta}_{\star})\|_{(\textbf{U}_{\star}^{\top}\textbf{V}_{t}\textbf{U}_{\star})^{-1}}$
$\leq RT\Delta_{\operatorname{off}}+\sum_{t=1}^{T}\alpha_{2,t}\|\phi(x_{t},a_{t})^{\top}\textbf{U}_{\star}\|_{(\textbf{U}_{\star}^{\top}\textbf{V}_{t}\textbf{U}_{\star})^{-1}}$
=$\stackrel{{\scriptstyle(ii)}}{{=}}RT\Delta_{\operatorname{off}}+\sum_{t=1}^{T}\alpha_{2,t}\|\phi(x_{t},a_{t})^{\top}\textbf{U}_{\star}\|_{\textbf{V}_{1,t}^{-1}}$
$\stackrel{{\scriptstyle(iii)}}{{\leq}}RT\Delta_{\operatorname{off}}+O\left(d_{K}\sqrt{T\log\left(T/\delta\right)}\right)$
=$\stackrel{{\scriptstyle(iv)}}{{=}}O\left(Rd_{K}\sqrt{T\log\left(T/\delta\right)}+\frac{R\lambda_{D}^{3}}{\lambda_{\min}}\sqrt{\frac{T^{2}d_{K}d_{A}\log(d_{A}/\delta)}{N}}\right)$
$=O\left(Rd_{K}\sqrt{T\log\left(T/\delta\right)}\left(1+\frac{\lambda_{D}^{3}}{\lambda_{\min}}\sqrt{\frac{Td_{A}\log(d_{A})}{d_{K}N}}\right)\right)$
$=\tilde{O}\left(Rd_{K}\sqrt{T}\left(1+\sqrt{\frac{Td_{A}}{d_{K}N}}\right)\right)$
where $(i)$ holds since $\tilde{\textbf{U}}_{t}\tilde{\textbf{U}}_{t}^{\top}\tilde{\bm{\beta}}_{t}=\tilde{\bm{\beta}}_{t}$ and $\textbf{U}_{\star}\textbf{U}_{\star}^{\top}\bm{\beta}_{\star}=\bm{\beta}_{\star}$ $(ii)$ holds since $\textbf{V}_{1,t}^{-1}=(\textbf{U}_{\star}^{\top}\textbf{V}_{t}\textbf{U}_{\star})^{-1}$, $(iii)$ holds by the usual proof of LinUCB applied to dimension $d_{K}$, and $(iv)$ holds by Theorem 2.
Leveraging optimism in high dimension: This is merely the proof of LinUCB. From above and by the optimistic design of the algorithm, with probability $1-\delta$, $\phi(x_{t},a_{t})^{\top}\tilde{\bm{\beta}}_{t}$ is an upper bound on $\phi(x_{t},a)^{\top}\bm{\beta}_{\star}$ for any action $a$. Thus, we have the following regret decomposition with probability at least $1-\delta$:
$\operatorname{Reg}_{T}$ $=\sum_{t=1}^{T}\phi(x_{t},a_{t}^{\star})^{\top}\bm{\beta}_{\star}-\phi(x_{t},a_{t})^{\top}\tilde{\bm{\beta}}_{\star}$
$\leq\sum_{t=1}^{T}\phi(x_{t},a_{t})^{\top}\tilde{\bm{\beta}}_{t}-\phi(x_{t},a_{t})^{\top}\bm{\beta}_{\star}$
$=\sum_{t=1}^{T}\|\phi(x_{t},a_{t})^{\top}\|_{\textbf{V}_{t}^{-1}}\|\tilde{\bm{\beta}}_{t}-\bm{\beta}_{\star}\|_{\textbf{V}_{t}}$
$=O(Rd_{A}\sqrt{T\log(T/\delta)})$
$=\tilde{O}(Rd_{A}\sqrt{T})$
where the last line follows from the standard regret bound for LinUCB applied to dimension $d_{A}$.
Combining the two bounds, we have our result.
$\operatorname{Reg}_{T}$ $=O\left(\min\left(Rd_{A}\sqrt{T\log(T/\delta)},Rd_{K}\sqrt{T\log\left(T/\delta\right)}\left(1+\frac{\lambda_{D}^{3}}{\lambda_{\min}}\sqrt{\frac{Td_{A}\log(d_{A})}{d_{K}N}}\right)\right)\right)$
$=\tilde{O}\left(\min\left(Rd_{A}\sqrt{T},Rd_{K}\sqrt{T}\left(1+\sqrt{\frac{Td_{A}}{d_{K}N}}\right)\right)\right)$
Appendix E Proofs for ProBALL-UCB
E.1 Confidence bound for the low-dimensional reward parameter
Lemma 6 (Confidence Bound for $\hat{\bm{\beta}}_{1,t}$).
If for all timesteps $t$, $\phi(x_{t},a_{t})$ lies in the span of $\hat{\textbf{U}}$, we have that for a universal constant $C$,
$\left\|\hat{\textbf{U}}^{\top}\hat{\bm{\beta}}_{1,t}-\hat{\textbf{U}}^{\top}\bm{\beta}_{\star}\right\|_{\textbf{V}_{1,t}^{-1}}\leq R\sqrt{\mu}+CR\sqrt{d_{A}\log(t/\delta)}$
Otherwise, we have that
$\left\|\hat{\textbf{U}}^{\top}\hat{\bm{\beta}}_{1,t}-\hat{\textbf{U}}^{\top}\bm{\beta}_{\star}\right\|_{\textbf{V}_{1,t}^{-1}}\leq R\sqrt{\mu}+R\kappa_{t}+\Delta_{\operatorname{off}}CR\sqrt{d_{A}\log(t/\delta)}$
where $\kappa_{t}=\|\hat{\textbf{U}}^{\top}\textbf{X}_{t}^{\top}\textbf{X}_{t}\|_{(\hat{\textbf{U}}^{\top}\textbf{V}_{t}\hat{\textbf{U}})^{-1}}$, with notation $\|\textbf{A}\|_{\textbf{C}}:=\sqrt{\|\textbf{A}^{\top}\textbf{C}\textbf{A}\|_{2}}$
Proof.
For brevity, in this proof we will denote by $\textbf{X}_{t}:=[\phi(x_{1},a_{1}),\dots\phi(x_{t-1},a_{t-1})]^{\top}$, which is a $t\times d_{A}$ matrix. Recall that $\textbf{V}_{t}=\mu\textbf{I}_{d_{A}}+\textbf{X}_{t}^{\top}\textbf{X}_{t}$. We will also denote $\textbf{V}_{1,t}=\hat{\textbf{U}}^{\top}\textbf{V}_{t}\hat{\textbf{U}}$. Note that the vector of rewards is given by $\textbf{X}_{t}\bm{\beta}_{\star}+\eta_{t}$, where $\eta_{t}$ is a random vector of $t$ independent $R^{2}$-subgaussian entries. So, $\textbf{b}_{t}=\textbf{X}_{t}^{\top}(\textbf{X}_{t}\bm{\beta}_{\star}+\eta_{t})$. Also define the notation $\Delta_{U}:=\hat{\textbf{U}}\hat{\textbf{U}}^{\top}-\textbf{U}_{\star}\textbf{U}_{\star}^{\top}$.
If all actions lie in the span of $\hat{\textbf{U}}$. Note that since all actions taken lie in the span of $\hat{\textbf{U}}$, $\textbf{X}_{t}=\textbf{X}_{t}\hat{\textbf{U}}\hat{\textbf{U}}^{\top}$. This is the key observation. Now note that
$\hat{\bm{\beta}}_{1,t}$ $=\hat{\textbf{U}}\textbf{V}_{1,t}^{-1}\hat{\textbf{U}}^{\top}\textbf{X}_{t}^{\top}(\textbf{X}_{t}\bm{\beta}_{\star}+\eta_{t})$
$=\hat{\textbf{U}}\textbf{V}_{1,t}^{-1}\hat{\textbf{U}}^{\top}\textbf{X}_{t}^{\top}\textbf{X}_{t}\bm{\beta}_{\star}+\hat{\textbf{U}}\textbf{V}_{1,t}^{-1}\hat{\textbf{U}}^{\top}\textbf{X}_{t}^{\top}\eta_{t}$
$=\hat{\textbf{U}}\textbf{V}_{1,t}^{-1}\hat{\textbf{U}}^{\top}\textbf{X}_{t}^{\top}\textbf{X}_{t}\hat{\textbf{U}}\hat{\textbf{U}}^{\top}\bm{\beta}_{\star}+\hat{\textbf{U}}\textbf{V}_{1,t}^{-1}\hat{\textbf{U}}^{\top}\textbf{X}_{t}^{\top}\eta_{t}$
$=\hat{\textbf{U}}\textbf{V}_{1,t}^{-1}(\textbf{V}_{1,t}-\mu\textbf{I}_{d_{K}})\hat{\textbf{U}}^{\top}\bm{\beta}_{\star}+\hat{\textbf{U}}\textbf{V}_{1,t}^{-1}\hat{\textbf{U}}^{\top}\textbf{X}_{t}^{\top}\eta_{t}$
$=\hat{\textbf{U}}\textbf{U}^{\top}\bm{\beta}_{\star}-\mu\hat{\textbf{U}}\textbf{V}_{1,t}^{-1}\hat{\textbf{U}}^{\top}\bm{\beta}_{\star}+\hat{\textbf{U}}\textbf{V}_{1,t}^{-1}\hat{\textbf{U}}^{\top}\textbf{X}_{t}^{\top}\eta_{t}$
The rest of the proof is similar to Theorem 2 in Abbasi-Yadkori et al. [2011]. We first note that for any $x$, we have that
$x^{\top}(\hat{\bm{\beta}}_{1,t}-\hat{\textbf{U}}\hat{\textbf{U}}^{\top}\bm{\beta}_{\star})$ $\leq\left|(x^{\top}\hat{\textbf{U}})\textbf{V}_{1,t}^{-1}(-\mu\hat{\textbf{U}}^{\top}\bm{\beta}_{\star}+\hat{\textbf{U}}^{\top}\textbf{X}_{t}^{\top}\eta_{t})\right|$
$\leq\|\hat{\textbf{U}}^{\top}x\|_{\textbf{V}_{1,t}^{-1}}\left\|(-\mu\hat{\textbf{U}}^{\top}\bm{\beta}_{\star}+\hat{\textbf{U}}^{\top}\textbf{X}_{t}^{\top}\eta_{t})\right\|_{\textbf{V}_{1,t}^{-1}}$
$\leq\|\hat{\textbf{U}}^{\top}x\|_{\textbf{V}_{1,t}^{-1}}\left(\mu\|\hat{\textbf{U}}^{\top}\bm{\beta}_{\star}\|_{\textbf{V}_{1,t}^{-1}}+\|\hat{\textbf{U}}^{\top}\textbf{X}_{t}^{\top}\eta_{t}\|_{\textbf{V}_{1,t}^{-1}}\right)$
Now note that $\mu\|\hat{\textbf{U}}^{\top}\bm{\beta}_{\star}\|_{\textbf{V}_{1,t}^{-1}}\leq\|\hat{\textbf{U}}^{\top}\bm{\beta}_{\star}\|_{2}\sqrt{\mu}\leq R\sqrt{\mu}$ and by the self normalized martingale concentration inequality from Abbasi-Yadkori et al. [2011] applied to $d_{K}$ dimensional vectors $\hat{\textbf{U}}^{\top}\textbf{X}_{t}$, we have that with probability at least $1-\delta$, $\|\hat{\textbf{U}}^{\top}\textbf{X}_{t}^{\top}\eta_{t}\|_{\textbf{V}_{1,t}^{-1}}\leq CR\sqrt{d_{A}\log(t/\delta)}$ for some universal constant $C$. So we have with probability at least $1-\delta$ that
$=\left\|\hat{\textbf{U}}(\hat{\bm{\beta}}_{1,t}-\hat{\textbf{U}}\hat{\textbf{U}}^{\top}\bm{\beta}_{\star})\right\|^{2}_{\textbf{V}_{1,t}^{-1}}$ $\leq\left\|\hat{\bm{\beta}}_{1,t}-\hat{\textbf{U}}\hat{\textbf{U}}^{\top}\bm{\beta}_{\star}\right\|_{\textbf{V}_{1,t}^{-1}}\left(R\sqrt{\mu}+CR\sqrt{d_{A}\log(t/\delta)}\right)$
This means that
$\left\|\hat{\textbf{U}}^{\top}\hat{\bm{\beta}}_{1,t}-\hat{\textbf{U}}^{\top}\bm{\beta}_{\star}\right\|_{\textbf{V}_{1,t}^{-1}}$ $=\left\|\hat{\bm{\beta}}_{1,t}-\hat{\textbf{U}}\hat{\textbf{U}}^{\top}\bm{\beta}_{\star}\right\|_{\textbf{V}_{1,t}^{-1}}$
$\leq R\sqrt{\mu}+CR\sqrt{d_{A}\log(t/\delta)}$
If all actions don’t lie in the span of $\hat{\textbf{U}}$. This time note that
$\hat{\bm{\beta}}_{1,t}$ $=\hat{\textbf{U}}\textbf{V}_{1,t}^{-1}\hat{\textbf{U}}^{\top}\textbf{X}_{t}^{\top}(\textbf{X}_{t}\bm{\beta}_{\star}+\eta_{t})$
$=\hat{\textbf{U}}\textbf{V}_{1,t}^{-1}\hat{\textbf{U}}^{\top}\textbf{X}_{t}^{\top}\textbf{X}_{t}\bm{\beta}_{\star}+\hat{\textbf{U}}\textbf{V}_{1,t}^{-1}\hat{\textbf{U}}^{\top}\textbf{X}_{t}^{\top}\eta_{t}$
$=\hat{\textbf{U}}\textbf{V}_{1,t}^{-1}\hat{\textbf{U}}^{\top}\textbf{X}_{t}^{\top}\textbf{X}_{t}\hat{\textbf{U}}_{\star}\textbf{U}_{\star}^{\top}\bm{\beta}_{\star}+\hat{\textbf{U}}\textbf{V}_{1,t}^{-1}\textbf{U}^{\top}\textbf{X}_{t}^{\top}\eta_{t}$
$=\hat{\textbf{U}}\textbf{V}_{1,t}^{-1}\hat{\textbf{U}}^{\top}\textbf{X}_{t}^{\top}\textbf{X}_{t}\hat{\textbf{U}}\hat{\textbf{U}}^{\top}\bm{\beta}_{\star}+\hat{\textbf{U}}\textbf{V}_{1,t}^{-1}\hat{\textbf{U}}^{\top}\textbf{X}_{t}^{\top}\textbf{X}_{t}\Delta_{U}\bm{\beta}_{\star}+\hat{\textbf{U}}\textbf{V}_{1,t}^{-1}\hat{\textbf{U}}^{\top}\textbf{X}_{t}^{\top}\eta_{t}$
$=\hat{\textbf{U}}\textbf{V}_{1,t}^{-1}(\textbf{V}_{1,t}-\mu\textbf{I}_{d_{K}})\hat{\textbf{U}}^{\top}\bm{\beta}_{\star}+\hat{\textbf{U}}\textbf{V}_{1,t}^{-1}\hat{\textbf{U}}^{\top}\textbf{X}_{t}^{\top}\textbf{X}_{t}\Delta_{U}\bm{\beta}_{\star}+\hat{\textbf{U}}\textbf{V}_{1,t}^{-1}\hat{\textbf{U}}^{\top}\textbf{X}_{t}^{\top}\eta_{t}$
$=\hat{\textbf{U}}\textbf{U}^{\top}\bm{\beta}_{\star}-\mu\hat{\textbf{U}}\textbf{V}_{1,t}^{-1}\hat{\textbf{U}}^{\top}\bm{\beta}_{\star}+\hat{\textbf{U}}\textbf{V}_{1,t}^{-1}\hat{\textbf{U}}^{\top}\textbf{X}_{t}^{\top}\textbf{X}_{t}\Delta_{U}\bm{\beta}_{\star}+\hat{\textbf{U}}\textbf{V}_{1,t}^{-1}\hat{\textbf{U}}^{\top}\textbf{X}_{t}^{\top}\eta_{t}$
The rest of the proof is similar to Theorem 2 in Abbasi-Yadkori et al. [2011]. We first note that for any $x$, we have that
$x^{\top}(\hat{\bm{\beta}}_{1,t}-\hat{\textbf{U}}\hat{\textbf{U}}^{\top}\bm{\beta}_{\star})$ $\leq\left|(x^{\top}\hat{\textbf{U}})\textbf{V}_{1,t}^{-1}(-\mu\hat{\textbf{U}}^{\top}\bm{\beta}_{\star}+\hat{\textbf{U}}^{\top}\textbf{X}_{t}^{\top}\textbf{X}_{t}\Delta_{U}\bm{\beta}_{\star}+\hat{\textbf{U}}^{\top}\textbf{X}_{t}^{\top}\eta_{t})\right|$
$\leq\|\hat{\textbf{U}}^{\top}x\|_{\textbf{V}_{1,t}^{-1}}\left\|(-\mu\hat{\textbf{U}}^{\top}\bm{\beta}_{\star}+\hat{\textbf{U}}^{\top}\textbf{X}_{t}^{\top}\textbf{X}_{t}\Delta_{U}\bm{\beta}_{\star}+\hat{\textbf{U}}^{\top}\textbf{X}_{t}^{\top}\eta_{t})\right\|_{\textbf{V}_{1,t}^{-1}}$
$\leq\|\hat{\textbf{U}}^{\top}x\|_{\textbf{V}_{1,t}^{-1}}\left(\mu\|\hat{\textbf{U}}^{\top}\bm{\beta}_{\star}\|_{\textbf{V}_{1,t}^{-1}}+\|\hat{\textbf{U}}^{\top}\textbf{X}_{t}^{\top}\textbf{X}_{t}\Delta_{U}\bm{\beta}_{\star}\|_{\textbf{V}_{1,t}^{-1}}+\|\hat{\textbf{U}}^{\top}\textbf{X}_{t}^{\top}\eta_{t}\|_{\textbf{V}_{1,t}^{-1}}\right)$
$\stackrel{{\scriptstyle(i)}}{{\leq}}\|\hat{\textbf{U}}^{\top}x\|_{\textbf{V}_{1,t}^{-1}}\left(\mu\|\hat{\textbf{U}}^{\top}\bm{\beta}_{\star}\|_{\textbf{V}_{1,t}^{-1}}+\|\Delta_{U}\bm{\beta}_{\star}\|_{2}\|\hat{\textbf{U}}^{\top}\textbf{X}_{t}^{\top}\textbf{X}_{t}\|_{\textbf{V}_{1,t}^{-1}}+\|\hat{\textbf{U}}^{\top}\textbf{X}_{t}^{\top}\eta_{t}\|_{\textbf{V}_{1,t}^{-1}}\right)$
$\leq\|\hat{\textbf{U}}^{\top}x\|_{\textbf{V}_{1,t}^{-1}}\left(\mu\|\hat{\textbf{U}}^{\top}\bm{\beta}_{\star}\|_{\textbf{V}_{1,t}^{-1}}+R\|\Delta_{U}\|_{2}\|\hat{\textbf{U}}^{\top}\textbf{X}_{t}^{\top}\textbf{X}_{t}\|_{\textbf{V}_{1,t}^{-1}}+\|\hat{\textbf{U}}^{\top}\textbf{X}_{t}^{\top}\eta_{t}\|_{\textbf{V}_{1,t}^{-1}}\right)$
$\leq\|\hat{\textbf{U}}^{\top}x\|_{\textbf{V}_{1,t}^{-1}}\left(\mu\|\hat{\textbf{U}}^{\top}\bm{\beta}_{\star}\|_{\textbf{V}_{1,t}^{-1}}+R\Delta_{\operatorname{off}}\|\hat{\textbf{U}}^{\top}\textbf{X}_{t}^{\top}\textbf{X}_{t}\|_{\textbf{V}_{1,t}^{-1}}+\|\hat{\textbf{U}}^{\top}\textbf{X}_{t}^{\top}\eta_{t}\|_{\textbf{V}_{1,t}^{-1}}\right)$
[…]
Now note that $\mu\|\hat{\textbf{U}}^{\top}\bm{\beta}_{\star}\|_{\textbf{V}_{1,t}^{-1}}\leq\|\hat{\textbf{U}}^{\top}\bm{\beta}_{\star}\|_{2}\sqrt{\mu}\leq R\sqrt{\mu}$ and by the self normalized martingale concentration inequality from Abbasi-Yadkori et al. [2011] applied to $d_{K}$ dimensional vectors $\hat{\textbf{U}}^{\top}\textbf{X}_{t}$, we have that with probability at least $1-\delta$, $\|\hat{\textbf{U}}^{\top}\textbf{X}_{t}^{\top}\eta_{t}\|_{\textbf{V}_{1,t}^{-1}}\leq CR\sqrt{d_{A}\log(t/\delta)}$ for some universal constant $C$. Also recall that $\|\hat{\textbf{U}}^{\top}\textbf{X}_{t}^{\top}\textbf{X}_{t}\|_{\textbf{V}_{1,t}^{-1}}=\kappa_{t}$. So we have with probability at least $1-\delta$ that
$\left\|\hat{\textbf{U}}(\hat{\bm{\beta}}_{1,t}-\hat{\textbf{U}}\hat{\textbf{U}}^{\top}\bm{\beta}_{\star})\right\|^{2}_{\textbf{V}_{1,t}^{-1}}$ $\leq\left\|\hat{\textbf{U}}^{\top}(\hat{\bm{\beta}}_{1,t}-\hat{\textbf{U}}\hat{\textbf{U}}^{\top}\bm{\beta}_{\star})\right\|_{\textbf{V}_{1,t}^{-1}}\left(R\sqrt{\mu}+R\Delta_{\operatorname{off}}\kappa_{t}+CR\sqrt{d_{A}\log(t/\delta)}\right)$
This means that
$\left\|\hat{\textbf{U}}^{\top}\hat{\bm{\beta}}_{1,t}-\hat{\textbf{U}}^{\top}\bm{\beta}_{\star}\right\|_{\textbf{V}_{1,t}^{-1}}$ $=\left\|\hat{\textbf{U}}(\hat{\bm{\beta}}_{1,t}-\hat{\textbf{U}}\hat{\textbf{U}}^{\top}\bm{\beta}_{\star})\right\|_{\textbf{V}_{1,t}^{-1}}$
$\leq R\sqrt{\mu}+R\Delta_{\operatorname{off}}\kappa_{t}+CR\sqrt{d_{A}\log(t/\delta)}$
Note that $\kappa_{t}=\left\|\sum_{s=1}^{t-1}\hat{\textbf{U}}^{\top}\phi(x_{s},a_{s})\phi(x_{s},a_{s})^{\top}\right\|_{\textbf{V}_{1,t}^{-1}}=\frac{1}{\sqrt{\mu}}\left\|\sum_{s=1}^{t-1}\hat{\textbf{U}}^{\top}\phi(x_{s},a_{s})\phi(x_{s},a_{s})^{\top}\right\|_{2}=O(t)$, but this is a worst case bound. ∎
We also state a lemma, borrowed from the standard proof of LinUCB regret.
Lemma 7.
For any sequence of actions and contexts $x_{t},a_{t}$ and $\textbf{V}_{t}=\mu\textbf{I}+\sum_{s=1}^{t-1}\phi(x_{t},a_{t})\phi(x_{t},a_{t})^{\top}$, we have that
$\sum_{t=1}^{T}\min\left(\|\phi(x_{t},a_{t})\|^{2}_{\textbf{V}_{t}^{-1}},1\right)=O(\sqrt{d_{A}})$
$\sum_{t=1}^{T}\min\left(\|\hat{\textbf{U}}^{\top}\phi(x_{t},a_{t})\|^{2}_{(\hat{\textbf{U}}^{\top}\textbf{V}_{t}\hat{\textbf{U}})^{-1}},1\right)=O(\sqrt{d_{K}})$
Proof.
This follows immediately from Lemma 11 of Abbasi-Yadkori et al. [2011]. ∎
E.2 Proof of the theorem
See 4
Proof.
We will first show that our bonuses give optimistic estimates of the true reward whp and then leverage the optimism.
Showing optimism when $\tau\Delta_{\operatorname{off}}\sqrt{T}\leq d_{A}$: In this case, we are inside the "if" statement and are running a projected and modified version of LinUCB. In that case, for all timesteps the projected version is run, all features will lie in the span of $\hat{\textbf{U}}$. This is because the maximization problem is given by
$a_{t}=\arg\max_{a,\|\phi(x_{t},a)\|_{2}\leq 1}\phi^{\top}\hat{\textbf{U}}\hat{\textbf{U}}^{\top}\hat{\bm{\beta}}_{1,t}+\|\hat{\textbf{U}}^{\top}\phi(x_{t},a)\|_{(\hat{\textbf{U}}^{\top}\textbf{V}_{1,t}\hat{\textbf{U}})^{-1}}$
If our features are isotropic, then this is only maximized by a feature in the span of $\hat{\textbf{U}}$. So, the conditions of the first bound in Lemma 6 are fulfilled. We will denote $\textbf{V}_{1,t}=\hat{\textbf{U}}^{\top}\textbf{V}_{t}\hat{\textbf{U}}$. We have that with probability $21-\delta/$ 2, for all $x,a,t$, the following holds.
$|\phi(x,a)^{\top}\bm{\beta}_{\star}-\phi(x,a)^{\top}\hat{\textbf{U}}\hat{\textbf{U}}^{\top}\hat{\bm{\beta}}_{1,t}|$ $\leq|\phi(x,a)^{\top}\hat{\textbf{U}}(\hat{\textbf{U}}^{\top}\bm{\beta}_{\star}-\hat{\textbf{U}}^{\top}\hat{\bm{\beta}}_{1,t})|+|\phi(x,a)^{\top}(\textbf{U}_{\star}\textbf{U}_{\star}^{\top}-\hat{\textbf{U}}\hat{\textbf{U}}^{\top})\bm{\beta}_{\star}|$
$\leq\|\hat{\textbf{U}}^{\top}\phi(x,a)\|_{\textbf{V}_{1,t}^{-1}}\|\hat{\textbf{U}}^{\top}\bm{\beta}_{\star}-\hat{\textbf{U}}^{\top}\hat{\bm{\beta}}_{1,t}\|_{\textbf{V}_{1,t}^{-1}}+R\|\textbf{U}_{\star}\textbf{U}_{\star}^{\top}-\hat{\textbf{U}}\hat{\textbf{U}}^{\top}\|_{2}$
$\leq\alpha_{1,t}\|\hat{\textbf{U}}^{\top}\phi(x,a)\|_{\textbf{V}_{1,t}^{-1}}+R\Delta_{\operatorname{off}}$
This shows that with probability at least $1-\delta$ and for all $x,a,t$,
$\phi(x,a)^{\top}\bm{\beta}_{\star}\leq\phi(x,a)^{\top}\hat{\textbf{U}}\hat{\textbf{U}}^{\top}\hat{\bm{\beta}}_{1,t}+\alpha_{1,t}\|\hat{\textbf{U}}^{\top}\phi(x,a)\|_{\textbf{V}_{1,t}^{-1}}+R\Delta_{\operatorname{off}}$
This implies that with probability at least $1-\delta$, for any action $a$,
$\phi(x_{t},a)^{\top}\bm{\beta}_{\star}$ $\leq\max_{a}\phi(x_{t},a)^{\top}\hat{\textbf{U}}\hat{\textbf{U}}^{\top}\hat{\bm{\beta}}_{1,t}+\alpha_{1,t}\|\hat{\textbf{U}}^{\top}\phi(x_{t},a)\|_{\textbf{V}_{1,t}^{-1}}+R\Delta_{\operatorname{off}}$
$=\phi(x_{t},a_{t})^{\top}\hat{\textbf{U}}\hat{\textbf{U}}^{\top}\hat{\bm{\beta}}_{1,t}+\alpha_{1,t}\|\hat{\textbf{U}}^{\top}\phi(x_{t},a_{t})\|_{\textbf{V}_{1,t}^{-1}}+R\Delta_{\operatorname{off}}$
Leveraging optimism when $\tau\Delta_{\operatorname{off}}\sqrt{T}\leq d_{A}$: Consider the standard regret decomposition, which holds with probability $1-\delta$:
$\operatorname{Reg}_{T}$ $=\sum_{t=1}^{T}\phi(x_{t},a_{t}^{\star})^{\top}\bm{\beta}_{\star}-\phi(x_{t},a_{t})^{\top}\bm{\beta}_{\star}$
$\leq\sum_{t=1}^{T}\phi(x_{t},a_{t})^{\top}\hat{\textbf{U}}\hat{\textbf{U}}^{\top}\hat{\bm{\beta}}_{1,t}-\phi(x_{t},a_{t})^{\top}\bm{\beta}_{\star}+\alpha_{1,t}\|\hat{\textbf{U}}^{\top}\phi(x_{t},a_{t})\|_{\textbf{V}_{1,t}^{-1}}+R\Delta_{\operatorname{off}}$
$\leq\sum_{t=1}^{T}2\alpha_{1,t}\|\hat{\textbf{U}}^{\top}\phi(x_{t},a_{t})\|_{\textbf{V}_{1,t}^{-1}}+2R\Delta_{\operatorname{off}}$ (2)
$\stackrel{{\scriptstyle(i)}}{{\leq}}\sum_{t=1}^{T}2\alpha^{\prime}_{1,t}\min\left(\|\hat{\textbf{U}}^{\top}\phi(x_{t},a_{t})\|_{\textbf{V}_{1,t}^{-1}},1\right)+2R\Delta_{\operatorname{off}}$
$\leq\sum_{t=1}^{T}2(R\sqrt{\mu}+C\sqrt{d_{K}\log(T/\delta)})\min\left(\|\hat{\textbf{U}}^{\top}\phi(x_{t},a_{t})\|_{\textbf{V}_{1,t}^{-1}},1\right)+2R\Delta_{\operatorname{off}}$
$\qquad+\sum_{t=1}^{T}2R\Delta_{\operatorname{off}}\kappa_{t}\min\left(\|\hat{\textbf{U}}^{\top}\phi(x_{t},a_{t})\|_{\textbf{V}_{1,t}^{-1}},1\right)$
$\stackrel{{\scriptstyle(ii)}}{{\leq}}2(R\sqrt{\mu}+C\sqrt{d_{K}\log(T/\delta)})\sqrt{\sum_{t=1}^{T}\min\left(\|\hat{\textbf{U}}^{\top}\phi(x_{t},a_{t})\|^{2}_{\textbf{V}_{1,t}^{-1}},1\right)}+2R\Delta_{\operatorname{off}}T$
$\qquad+2R\Delta_{\operatorname{off}}\sqrt{\sum_{t=1}^{T}\kappa^{2}_{t}}\sqrt{\sum_{t=1}^{T}\min\left(\|\hat{\textbf{U}}^{\top}\phi(x_{t},a_{t})\|^{2}_{\textbf{V}_{1,t}^{-1}},1\right)}$
$\stackrel{{\scriptstyle(iii)}}{{\leq}}O(d_{K}\sqrt{T\log(T/\delta)})+2R\Delta_{\operatorname{off}}T+2R\Delta_{\operatorname{off}}\sqrt{\frac{\sum_{t=1}^{T}\kappa^{2}_{t}}{T}}\sqrt{d_{K}T}$ (3)
$=O\left(Rd_{K}\sqrt{T\log(T/\delta)}\left(1+\Delta_{\operatorname{off}}\left(\frac{\sqrt{T}}{d_{K}}+\sqrt{\frac{1}{T{d_{K}}}\sum_{t=1}^{T}\kappa_{t}^{2}}\right)\right)\right)$
where $(i)$ holds since $\phi(x_{t},a_{t}^{\star})^{\top}\bm{\beta}_{\star}-\phi(x_{t},a_{t})^{\top}\bm{\beta}_{\star}\leq 2R$, $(ii)$ holds by the Cauchy Schwarz inequality and $(iii)$ holds by Lemma 7. So, we have that
$\operatorname{Reg}_{T}$ $=O\left(Rd_{K}\sqrt{T\log\left(T/\delta\right)}\left(1+\frac{1}{\lambda_{A}^{3}\lambda_{\theta}}\sqrt{\frac{d_{A}\log(d_{A})}{Nd_{K}}}\left(\sqrt{{T}}+\sqrt{\frac{d_{K}}{T}\sum_{t=1}^{S}\kappa_{t}^{2}}\right)\right)\right)$
$=\widetilde{O}\left(Rd_{K}\sqrt{T}\left(1+\frac{1}{\lambda_{A}^{3}\lambda_{\theta}}\left(\sqrt{\frac{d_{A}T}{Nd_{K}}}+\sqrt{\frac{d_{A}}{TN}\sum_{t=1}^{T}\kappa_{t}^{2}}\right)\right)\right)$
$=\widetilde{O}\left(Rd_{K}\sqrt{T}\left(1+\frac{1}{\lambda_{A}^{3}\lambda_{\theta}}\left(\sqrt{\frac{d_{A}T}{Nd_{K}}}+\sqrt{\frac{d_{A}}{\min(S,T)N}\sum_{t=1}^{\min(S,T)}\kappa_{t}^{2}}\right)\right)\right)$
where the last inequality holds since $T<S$ for the first timestep $S$ satisfying $\Delta_{\operatorname{off}}\left(\sqrt{S}+\sqrt{\frac{d_{K}}{S}\sum_{t=1}^{S}\kappa_{t}^{2}}\right)\geq d_{A}$. Additionally, since $\Delta_{\operatorname{off}}\left(\sqrt{T}+\sqrt{\frac{d_{K}}{T}\sum_{t=1}^{T}\kappa_{t}^{2}}\right)\leq d_{A}$, we have using equation 3 that
$\operatorname{Reg}_{T}=\widetilde{O}(d_{A}\sqrt{T})$
as well. So, we have that when $\tau\Delta_{\operatorname{off}}\sqrt{T}\leq d_{A}$,
$\operatorname{Reg}_{T}$ $=\widetilde{O}\left(\min\left(d_{A}\sqrt{T},Rd_{K}\sqrt{T}\left(1+\frac{1}{\lambda_{A}^{3}\lambda_{\theta}}\left(\sqrt{\frac{d_{A}T}{Nd_{K}}}+\sqrt{\frac{d_{A}}{\min(S,T)N}\sum_{t=1}^{\min(S,T)}\kappa_{t}^{2}}\right)\right)\right)\right)$
Bounding regret when $\tau\Delta_{\operatorname{off}}\left(\sqrt{T}+\sqrt{\frac{d_{K}}{T}\sum_{t=1}^{T}\kappa_{t}^{2}}\right)\geq d_{A}$: In this regime, after the first timestep $S$ satisfying $\Delta_{\operatorname{off}}\left(\sqrt{S}+\sqrt{\frac{d_{K}}{S}\sum_{t=1}^{S}\kappa_{t}^{2}}\right)\geq d_{A}$, we run standard LinUCB with dimension $d_{A}$ and incur $\widetilde{O}(d_{A}\sqrt{T})$ regret with probability $1-\delta$. Until timestep $S-1$, we run the projected and modified version of LinUCB and incur regret bounded by
$\widetilde{O}(d_{K}\sqrt{S})+2R\Delta_{\operatorname{off}}S+2R\Delta_{\operatorname{off}}\sqrt{\frac{\sum_{t=1}^{S}\kappa^{2}_{t}}{S}}\sqrt{d_{K}S}$ $=\widetilde{O}\left(d_{A}\sqrt{S}+d_{K}\sqrt{S}\right)$
$=\widetilde{O}(d_{A}\sqrt{T}+d_{K}\sqrt{T})$
$=\widetilde{O}(d_{A}\sqrt{T})$
Combining these, we incur $O(d_{A}\sqrt{T})$ regret during the whole method.
Also, since $d_{A}\sqrt{T}\leq\Delta_{\operatorname{off}}\left(\sqrt{TS}+\sqrt{\frac{d_{K}T}{S}\sum_{t=1}^{S}\kappa_{t}^{2}}\right)$, we get that our regret is also bounded by
$\operatorname{Reg}_{T}$ $=\widetilde{O}\left(\Delta_{\operatorname{off}}\left(\sqrt{TS}+\sqrt{\frac{d_{K}T}{S}\sum_{t=1}^{S}\kappa_{t}^{2}}\right)\right)$
$=\widetilde{O}(d_{K}\sqrt{T}+\Delta_{\operatorname{off}}\left(\sqrt{TS}+\sqrt{\frac{d_{K}T}{S}\sum_{t=1}^{S}\kappa_{t}^{2}}\right))$
$=\widetilde{O}\left(Rd_{K}\sqrt{T}\left(1+\frac{1}{\lambda_{A}^{3}\lambda_{\theta}}\left(\sqrt{\frac{d_{A}T}{Nd_{K}}}+\sqrt{\frac{d_{A}}{\min(S,T)N}\sum_{t=1}^{\min(S,T)}\kappa_{t}^{2}}\right)\right)\right)$
So, our regret satisfies
$\operatorname{Reg}_{T}$ $=\widetilde{O}\left(\min\left(d_{A}\sqrt{T},Rd_{K}\sqrt{T}\left(1+\frac{1}{\lambda_{A}^{3}\lambda_{\theta}}\left(\sqrt{\frac{d_{A}T}{Nd_{K}}}+\sqrt{\frac{d_{A}}{\min(S,T)N}\sum_{t=1}^{\min(S,T)}\kappa_{t}^{2}}\right)\right)\right)\right)$
E.2.1 Understanding $\kappa_{t}$
Letting $\textbf{X}_{t}=\hat{\textbf{U}}^{\top}[\phi(x_{1},a_{1}),\dots\phi(x_{t},a_{t})]^{\top}$, recall that $\kappa_{t}:=\|\hat{\textbf{U}}\textbf{X}_{t}^{\top}\textbf{X}_{t}\|_{(\mu\textbf{I}+\textbf{X}_{t}^{\top}\textbf{X}_{t})^{-1}}$. In the worst case, $\kappa_{t}\leq\|\hat{\textbf{U}}\textbf{X}_{t}^{\top}\textbf{X}_{t}\|_{2}=O(t)$ since actions have norm $1$. In that case, $\frac{1}{S}\sum_{t=1}^{S}\kappa_{t}=O(S^{2})$. However, if $\textbf{X}_{t}$ lies in the span of $\hat{\textbf{U}}^{\top}$, then $\textbf{X}_{t}\hat{\textbf{U}}^{\top}\hat{\textbf{U}}=\textbf{X}_{t}$. This means that for $\textbf{Y}_{t}:=\textbf{X}_{t}\hat{\textbf{U}}^{\top}$,
$\kappa_{t}$ $=\sqrt{\hat{\textbf{U}}\textbf{X}_{t}^{\top}\textbf{X}_{t}\hat{\textbf{U}}^{\top}\hat{\textbf{U}}(\mu\textbf{I}+\textbf{X}_{t}^{\top}\textbf{X}_{t})^{-1}\textbf{U}^{\top}\hat{\textbf{U}}\textbf{X}_{t}^{\top}\textbf{X}_{t}\hat{\textbf{U}}^{\top}}$
$=\sqrt{\textbf{Y}_{t}^{\top}\textbf{Y}_{t}(\mu I+\textbf{Y}_{t}^{\top}\textbf{Y}_{t})^{-1}\textbf{Y}_{t}^{\top}\textbf{Y}_{t}}$
$=\sqrt{(\textbf{I}-\mu(\mu\textbf{I}+\textbf{Y}_{t}^{\top}\textbf{Y}_{t}))\textbf{Y}_{t}^{\top}\textbf{Y}_{t}}$
$=\widetilde{O}(\sqrt{t})$
So, $\frac{1}{S}\sum_{t=1}^{S}\kappa_{t}=O(S)=O(T)$.
Appendix F Other Experiments
F.1 Determining the Latent Rank from Offline Data
We note that as discussed in 4, we can use the eigenvalues of $\overline{\textbf{D}}_{N,1}^{-1}{\overline{\textbf{M}}_{N}}{\overline{\textbf{D}}_{N,2}^{-1}}$ to determine the rank of our subspace. We use the version of this arising from pseudo-inverses instead of regularization, just like in the MovieLens experiments. We demonstrate that we can indeed determine that the $d_{K}=18$ by finding the significant eigenvalues of the pseudo-inverse version of $\overline{\textbf{D}}_{N,1}^{-1}{\overline{\textbf{M}}_{N}}{\overline{\textbf{D}}_{N,2}^{-1}}$ estimated from the offline dataset of 5000 samples. We show the plots and log plots of these eigenvalues. We also plot the eigenvalues of the completed ratings matrix for comparison. Notice that they match and both fall after $18$ eigenvalues.
Figure 4: Plot of eigenvalues of aforementioned matrix. Notice the drop after $18$ eigenvalues.
Figure 5: Log-plot of eigenvalues of aforementioned matrix. Notice the drop after $18$ eigenvalues.
Appendix G Additional Figures
Figure 6: Comparison of ProBALL-UCB initialized with ground truth subspaces against LinUCB, mUCB, and mmUCB, for different choices of $\tau$ and confidence bound constructions. All variants of ProBALL-UCB perform no worse than LinUCB, and outperform mUCB and mmUCB. Shaded area depicts $1$-standard error confidence intervals over 30 trials with fresh $\theta$. The confidence intervals on regret thus account for the variation in frequentist regret for changing $\theta$.
Figure 7: Comparison of ProBALL-UCB initialized with ground truth subspaces against LinUCB, mUCB, and mmUCB, for different choices of $\tau$ and confidence bound constructions, in terms of rolling average rating over $25$ timesteps. All variants of ProBALL-UCB perform no worse than LinUCB, and outperform mUCB and mmUCB. Shaded area depicts $1$-standard error confidence intervals over 30 trials with fresh $\theta$. The confidence intervals on regret thus account for the variation in frequentist regret for changing $\theta$.
Figure 8: Comparison of ProBALL-UCB initialized with SOLD against LinUCB, mUCB, and mmUCB, for different choices of $\tau$ and confidence bound constructions, in terms of rolling average rating over $25$ timesteps. All variants of ProBALL-UCB perform no worse than LinUCB, and outperform mUCB and mmUCB. Shaded area depicts $1$-standard error confidence intervals over 30 trials with fresh $\theta$. The confidence intervals on regret thus account for the variation in frequentist regret for changing $\theta$.